### v2

# The Tension of Space as Dark Energy: A No Geometric Sequestering Theorem for Minimal Sequesters

Abstract
We model space as an elastic membrane and identify its uniform tension $T_{s}$ with the vacuum energy density of space. The central result is a no geometric sequestering theorem. In any minimal matter vacuum sequester framework where a global constraint cancels constant contributions from the matter sector by adjusting a global variable, a purely geometric unit operator remains intact. Concretely, a volume term of the form $S_{s}=-T_{s}\int\sqrt{-g}\,d^{4}x$ does not couple to the global variable that enforces the matter cancellation, and it survives unchanged in the local field equations. We establish this result within a covariant effective field theory using the background field method. The analysis shows that the coefficient $T_{s}$ is renormalized by graviton loops and in general by mixed matter–graviton loops and therefore runs with the renormalization scale $\mu$. The running defines an effective cosmological term $T_{s}(\mu)$ that is present even after the matter vacuum energy has been neutralized by the sequester constraint. In particular, we find that $T_{s}(\mu)$ sources $G_{\mu\nu}=-T_{s}(\mu)\,g_{\mu\nu}+\cdots$. The tension term $T_{s}(\mu)$ can then be understood as residual dark energy in the minimal sequestering framework.
1 Introduction
When Einstein wrote the field equations he and many contemporaries believed the universe was static. A closer inspection showed that with homogeneous nonzero energy density a static solution does not exist for the original equations. In 1917 he added a cosmological term $\Lambda g_{\mu\nu}$ [2] to allow a static closed universe, which leads to
$R_{\mu\nu}-\frac{1}{2}R\,g_{\mu\nu}+\Lambda g_{\mu\nu}=\frac{8\pi G}{c^{4}}T_{\mu\nu}$ (1.1)
With this term it was clear that a homogeneous and isotropic closed universe with nonzero energy density $\rho$ can be static if $\Lambda=1/a^{2}=4\pi G\rho/c^{2}$, where $a$ is the curvature radius. After Hubble’s discovery of cosmic expansion [3] and the theoretical developments of Friedmann [4] and Lemaître [5], Einstein abandoned the cosmological term, reportedly calling it his biggest blunder.
Late in the 1990s the situation changed. Two supernova projects, the Supernova Cosmology Project [7] and the High-Z Supernova Search Team [6], measured Type Ia supernovae as standardizable candles [^1] Type Ia supernovae were used as standard candles due to their remarkably uniform light curves. and found them to be dimmer than expected in a decelerating model. The inferred expansion is accelerating. Independent evidence from the Cosmic Microwave Background anisotropies by DASI [8], WMAP [9], and Planck [10], and from Baryon Acoustic Oscillations [11, 12], supports this conclusion. For a classic review see [13].
Vacuum Energy and the Cosmological Term
The cosmological term invites an interpretation as vacuum energy. The vacuum should look the same to all inertial observers, so any stress it contributes must be Lorentz invariant in locally inertial frames. The only symmetric rank two tensor with this property is $\eta_{\mu\nu}$, which implies,
$T^{(\mathrm{vac})}_{\mu\nu}=-\rho_{\mathrm{vac}}\,c^{2}\,\eta_{\mu\nu}$ (1.2)
In arbitrary coordinates this becomes,
$T^{(\mathrm{vac})}_{\mu\nu}=-\rho_{\mathrm{vac}}\,c^{2}\,g_{\mu\nu}$ (1.3)
with $\rho_{\mathrm{vac}}\,c^{2}$ constant in spacetime (to preserve Lorentz invariance [14]).
The same form follows from a perfect fluid with equation of state $p=\omega\rho c^{2}$. The fluid stress tensor is [14],
$T_{\mu\nu}=\left(\rho+\frac{p}{c^{2}}\right)\,u_{\mu}\,u_{\nu}+p\,g_{\mu\nu}$ (1.4)
Using a barotropic equation of state $\omega=-1$, we have $p=-\rho_{\mathrm{vac}}\,c^{2}$ and therefore we obtain for the vacuum,
$T^{(\mathrm{vac})}_{\mu\nu}=\left(\rho_{vac}-\rho_{vac}\right)\,u_{\mu}\,u_{\nu}-{\rho_{vac}\ c^{2}}g_{\mu\nu}=-\rho_{\mathrm{vac}}\,c^{2}\,g_{\mu\nu}$ (1.5)
Thus a vacuum with constant energy density acts gravitationally like a perfect fluid with $\omega=-1$. It is natural to split the total stress tensor into matter plus radiation and vacuum parts,
$T_{\mu\nu}=T^{(M+R)}_{\mu\nu}+T^{(\mathrm{vac})}_{\mu\nu}$ (1.6)
Inserting this into Einstein’s original equations without $\Lambda$, we get,
$R_{\mu\nu}-\frac{1}{2}R\,g_{\mu\nu}=\frac{8\pi G}{c^{4}}\left(T^{(M+R)}_{\mu\nu}+T^{(\mathrm{vac})}_{\mu\nu}\right)$ (1.7)
and using $T^{(\mathrm{vac})}_{\mu\nu}=-\rho_{\mathrm{vac}}\,c^{2}\,g_{\mu\nu}$, we obtain,
$R_{\mu\nu}-\frac{1}{2}R\,g_{\mu\nu}=\frac{8\pi G}{c^{4}}T^{(M+R)}_{\mu\nu}-\frac{8\pi G\rho_{\mathrm{vac}}}{c^{2}}g_{\mu\nu}$ (1.8)
Defining $\Lambda\equiv 8\pi G\rho_{\mathrm{vac}}/c^{2}$, one recovers the equations with a cosmological term,
$R_{\mu\nu}-\frac{1}{2}R\,g_{\mu\nu}+\Lambda g_{\mu\nu}=\frac{8\pi G}{c^{4}}T^{(M+R)}_{\mu\nu}$ (1.9)
Hence the cosmological term is equivalent to vacuum energy in the field equations. Einstein introduced $\Lambda$ for a static model, but the same term arises from the presence of vacuum energy even without that motivation.
Dark Energy and the Cosmological Constant Problem
The vacuum contribution that drives late time acceleration is often called dark energy. It is dark because it does not radiate or absorb in a way we can directly detect, and its evidence comes from dynamics such as the acceleration rate. The measured mean energy density today is about $5.249\times 10^{-10}\,\mathrm{J\,m^{-3}}$. A naive quantum field theory estimate of the zero point energy gives a value larger by about $122$ orders of magnitude. This huge discrepancy is the cosmological constant problem. The question it raises is simple to state but hard to answer. What sets the tiny value of the vacuum energy that gravitates on cosmic scales, if the quantum contributions of known fields appear so large in simple estimates?
Many ideas have been explored. Representative proposals include quintessence [15, 16], k-essence [17, 18], modified gravity [19, 20, 21, 22, 23], the Dvali Gabadadze Porrati braneworld [24, 25], Chaplygin gas [26], holographic dark energy models [27, 28, 29], quantum cosmology [54], dynamical relaxation of the Cosmological Constant via membrane nucleation events of top forms [57] and multiple top forms [58]. For broad reviews see [30, 31].
Our goal here is to give a simple and intuitive origin for the cosmological term, and hence for dark energy, using only basic ideas that are compatible with but not restricted to string theory. The key input is the physics of extended objects with tension. We use this as a guide to show that a stress which is proportional to the metric arises naturally when a three dimensional medium with a constant volume tension fills space. We then argue that ordinary space itself can be modeled as such a medium. The resulting stress tensor has the same form as vacuum energy.
We begin with two parallel statements. For an infinitely extended fundamental three brane one finds,
$T_{\mu\nu(3)}=-T_{3}\,g_{\mu\nu}$ (1.10)
where $T_{3}$ is the brane tension with dimensions of energy density. If instead we treat space itself as an elastic three brane with an intrinsic volume tension $T_{s}$, then,
$T_{\mu\nu(s)}=-T_{s}\,g_{\mu\nu}$ (1.11)
Both expressions have the same structure as the vacuum stress tensor discussed earlier. The second statement is stronger conceptually because it identifies space itself as the source of the vacuum energy. We will analyze how a geometric tension of space and a four form sequestering structure interact with the vacuum sector and we will show that while matter vacuum energy can be neutralized by global constraints, a purely geometric NG-like tension cannot be removed. The residual cosmological term is then a geometric constant that we denote by $T_{s}$ and we identify it with Dark Energy.
2 Modeling Space as an Elastic Membrane
2.1 Review of Nambu–Goto Dynamics for Strings and Branes
A fundamental string in $1+1$ dimensions has worldsheet action,
$S_{1}=-T_{1}\int d^{2}\xi\,\sqrt{-\det(\gamma_{\alpha\beta})}$ (2.1)
with worldsheet coordinates $\xi^{\alpha}=(\tau,\sigma)$, string tension $T_{1}$ and induced metric,
$\gamma_{\alpha\beta}=\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}g_{\mu\nu}$ (2.2)
For a $d$ brane in (d+1)(d{+}1) dimensions the natural generalization is [32, 33, 34],
$S_{d}=-T_{d}\int d^{d+1}\xi\,\sqrt{(-1)^{d}\det(\gamma_{\alpha\beta})}$ (2.3)
where $T_{d}$ has dimensions of energy density in $d$ spatial dimensions. For $d\geq 3$ it is common to call $T_{d}$ a volume tension, since it generalizes surface tension to higher codimension.
We will specifically focus on the $d=3$ case,
$S_{3}=-T_{3}\int d^{4}\xi\,\sqrt{-\det(\gamma_{\alpha\beta})}$ (2.4)
2.2 Stress Tensor of an Infinitely Extended Three Brane
The stress tensor that couples to the ambient metric is defined by,
$T^{\mu\nu}_{(3)}(x)=\frac{2}{\sqrt{-g(x)}}\,\frac{\delta S_{3}}{\delta g_{\mu\nu}(x)}$ (2.5)
To vary $S_{3}$, note the standard determinant identity,
$\delta\sqrt{-\det\gamma}=\frac{1}{2}\sqrt{-\det\gamma}\,\gamma^{\alpha\beta}\delta\gamma_{\alpha\beta}$ (2.6)
Using $\delta\gamma_{\alpha\beta}=\partial_{\alpha}X^{\mu}\partial_{\beta}X^{\nu}\,\delta g_{\mu\nu}$, one finds,
$\frac{\delta S_{3}}{\delta g_{\mu\nu}(x)}=-\frac{T_{3}}{2}\int d^{4}\xi\,\sqrt{-\det\gamma}\,\gamma^{\alpha\beta}\,\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}\,\delta^{(4)}\big(x-X(\xi)\big)$ (2.7)
Thus,
$T^{\mu\nu}_{(3)}(x)=-\frac{T_{3}}{\sqrt{-g(x)}}\int d^{4}\xi\,\sqrt{-\det\gamma}\,\gamma^{\alpha\beta}\,\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}\,\delta^{(4)}\big(x-X(\xi)\big)$ (2.8)
For an infinitely extended brane the natural gauge choice is the static gauge. By choosing this gauge, we are basically identifying the time coordinate $\tau$ and the spatial coordinates $\sigma^{1},\sigma^{2},\sigma^{3}$ on the brane’s worldsheet with the coordinate time $t$ and spatial coordinates $x,y$ and $z$ of the target spacetime. In essence, the spacetime embedding coordinates of the brane simply become $X^{0}=\tau$ and $X^{i}=\sigma^{i}$ for $i=1,2,3$. Therefore, in the static gauge, we have,
$\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}=\delta^{\mu}_{\alpha}$ (2.9)
where $\delta^{\mu}_{\alpha}$ is the Kronecker delta. Furthermore, equation (2.2) dictates that the induced metric $\gamma_{\alpha\beta}$ on the brane world-volume is related to the target spacetime metric $g_{\mu\nu}$ as,
$\gamma_{\alpha\beta}=g_{\mu\nu}\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}$ (2.10)
Substituting (2.9) in the equation above, we get,
$\gamma_{\alpha\beta}=g_{\mu\nu}\ \delta^{\mu}_{\alpha}\ \delta^{\nu}_{\beta}$ (2.11)
which simply reduces to,
$\gamma_{\alpha\beta}=g_{\alpha\beta}$ (2.12)
Therefore, we see that the induced metric $\gamma_{\alpha\beta}$ on the $3$-brane world-volume is equal to the target spacetime metric $g_{\alpha\beta}$. Furthermore, the equality above directly implies that $\sqrt{-\det(\gamma_{\alpha\beta})}=\sqrt{-g}$ and $\gamma^{\alpha\beta}=g^{\alpha\beta}$. This simplifies the expression for the $3$-brane stress energy tensor $T^{\mu\nu}_{(3)}$ (equation (2.8)) to,
$T^{\mu\nu}_{(3)}=-T_{3}\int d^{4}\xi\ g^{\alpha\beta}\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}\delta^{(3+1)}(x-X^{\mu}(\xi))$ (2.13)
In order to evaluate the expression above, we can re-write it as,
$T^{\mu\nu}_{(3)}=-T_{3}\int_{-\infty}^{\infty}d\tau\int_{-\infty}^{\infty}d^{3}\sigma\ g^{\alpha\beta}\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}\delta(x^{0}-\tau)\prod_{1}^{3}\delta(x^{i}-\sigma^{i})$ (2.14)
which allows us to straightforwardly perform the integral with respect to $\tau$ and obtain,
$T^{\mu\nu}_{(3)}=-T_{3}\int_{-\infty}^{\infty}d^{3}\sigma\ g^{\alpha\beta}\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}\prod_{1}^{3}\delta(x^{i}-\sigma^{i})$ (2.15)
Furthermore, given that we have assumed that the brane is infinitely large, we have,
$\int_{-\infty}^{\infty}d^{3}\sigma\prod_{1}^{3}\delta(x^{i}-\sigma^{i})=1$
since $x^{i}$ is always within the range of $\sigma^{i}$ $\forall\ i$. Therefore we get,
$T^{\mu\nu}_{(3)}=-T_{3}\ g^{\alpha\beta}\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}$ (2.16)
Using equation (2.9) in the equation above, we obtain,
$T^{\mu\nu}_{(3)}=-T_{3}\ g^{\alpha\beta}\ \delta^{\mu}_{\alpha}\ \delta^{\nu}_{\beta}$ (2.17)
which gives,
$T^{\mu\nu}_{(3)}=-T_{3}\ g^{\mu\nu}$ (2.18)
and finally, we obtain,
$T_{\mu\nu(3)}=-T_{3}\ g_{\mu\nu}$ (2.19)
as claimed in Sec. 1. This matches the expected units because $T_{3}$ is an energy density. Comparing with equation (1.5) in the Introduction we see that a uniform three brane tension $T_{3}$ behaves gravitationally like a vacuum energy density $\rho_{\mathrm{vac}}c^{2}$.
2.3 Space as the Three Brane
The previous derivation invites a sharper hypothesis. Instead of positing a separate embedded brane, we may treat space itself as a three dimensional elastic medium with an intrinsic volume tension $T_{s}$. The worldvolume of space is the physical spacetime and its worldvolume metric is simply $g_{\mu\nu}$. The Nambu–Goto like action for space is then,
$S_{s}=-T_{s}\int d^{4}x\,\sqrt{-g}$ (2.20)
This has the same dimensions as equation (2.4) with $T_{3}$ replaced by $T_{s}$. It also perfectly mirrors the cosmological piece of the Einstein–Hilbert action. The latter reads,
$S_{\mathrm{EH}}=\int d^{4}x\sqrt{-g}\left(\frac{c^{4}}{16\pi G}\,(R-2\Lambda)+\mathcal{L}_{M}\right)$ (2.21)
in which the cosmological piece is,
$S_{\Lambda}=-\frac{\Lambda c^{4}}{8\pi G}\int d^{4}x\sqrt{-g}$ (2.22)
Using $\Lambda=8\pi G\,\rho_{\mathrm{vac}}/c^{2}$ in the equation above, we get,
$S_{\Lambda}=-\rho_{\mathrm{vac}}c^{2}\int d^{4}x\sqrt{-g}$ (2.23)
Hence, dimensionally $\rho_{\mathrm{vac}}c^{2}]=[T_{s}]=[T_{3}]$, and $S_{s}$ given by equation (2.20) exactly mimics $S_{\Lambda}$ at the level of local dynamics.
We now look at the stress tensor that follows from the elastic action of space equation (2.20). On varying $S_{s}$, we obtain,
$\delta S_{s}=-T_{s}\int d^{4}x\,\delta\sqrt{-g}=-\frac{T_{s}}{2}\int d^{4}x\,\sqrt{-g}\,g^{\mu\nu}\delta g_{\mu\nu}$ (2.24)
Therefore,
$\frac{2}{\sqrt{-g}}\,\frac{\delta S_{s}}{\delta g_{\mu\nu}}=T_{s}\,g^{\mu\,\nu}$ (2.25)
But, the left hand side is the definition of the stress tensor for space, that is,
$T^{\mu\nu}_{(s)}=\frac{2}{\sqrt{-g}}\,\frac{\delta S_{s}}{\delta g_{\mu\nu}}$ (2.26)
Therefore, we have,
$T^{\mu\nu}_{(s)}=-T_{s}\,g^{\mu\nu}$ (2.27)
or that,
$T_{\mu\nu(s)}=-T_{s}\,g_{\mu\nu}$ (2.28)
Thus the elastic model of space reproduces the vacuum stress tensor with the identification $T_{s}=\rho_{\mathrm{vac}}c^{2}$.
Therefore, the Nambu–Goto structure ensures that a uniform volume tension produces a stress proportional to the metric. For an infinitely extended three brane this gives $T_{\mu\nu(3)}=T_{3}g_{\mu\nu}$. When we elevate space itself to the role of the three brane with intrinsic volume tension $T_{s}$, the same computation yields $T_{\mu\nu(s)}=-T_{s}g_{\mu\nu}$. Hence the cosmological term can be interpreted as the elastic volume tension of space. In the next section we will couple this picture to a four form sequestering framework and show that while matter vacuum energy can be neutralized by global constraints, a geometric volume tension cannot be removed. The residual dark energy is then precisely this $T_{s}$.
It is important to mention here that we do not assume that physical space is embedded in a higher dimensional ambient manifold. The elastic membrane language in this paper is an intrinsic description and all observables are built from the intrinsic metric and its derived tensors. Introducing an embedding is possible as a change of variables but it adds redundant structure that is not needed for the dynamics we study. An embedding also introduces an arbitrary ambient metric that does not enter any measurement performed by observers confined to the physical spacetime. Our equations of motion are fully determined by intrinsic geometry, so the theory remains sensible without any reference to extrinsic data. For these reasons we avoid committing to an embedding and keep the framework intrinsic from the outset.
3 A No-Geometric-Neutralization Theorem for Sequestering Four–Forms
We now work in natural units $\hbar=c=1$ in $3+1$ dimensions and introduce the sequester action which includes the usual Einstein–Hilbert term, a fundamental $3$-form Nambu-Goto action for space (essentially treating it as an elastic membrane), a matter sector that couples to a globally scaled metric through a single global variable $\lambda$ and two four-form sectors $F_{4}$ and $\hat{F}_{4}$ that enforce global constraints, see [61, 62, 63, 64, 65] for the detailed framework. The action has the following form,
$S=$ $\int d^{4}x\,\sqrt{-g}\,\Big(\frac{M_{*}^{2}}{2}\,R-T_{s}-\Lambda\Big)+\int d^{4}x\,\sqrt{-g}\,L_{m}(\lambda^{-2}\,g_{\mu\nu},\psi)$
$+\sigma\Big(\frac{\Lambda}{\mu^{4}}\Big)\,\int F_{4}+\hat{\sigma}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\,\int\hat{F}_{4}$ (3.1)
with $M_{*}^{2}=1/8\pi G$. Note that we are working in a convention where the units of $\Lambda$ are $\Lambda]=\rm mass^{4}$ (energy density). The matter fields are collected into $\psi$ and the scale $\mu$ is a fixed reference scale. A note on notation, content of the action and minimal assumptions are in order,
a) We work with a Lorentzian metric $g_{\mu\nu}$ of signature (+++)(-+++). The Levi–Civita connection is metric compatible and curvature tensors have the usual definitions.
b) The total action includes the following pieces:
$\bullet$ Gravity: $\int\sqrt{-g}\,\frac{M_{*}^{2}}{2}\,R$.
$\bullet$ Geometric tension of space: $-\,T_{s}\int\sqrt{-g}$. One may think of $T_{s}$ as a constant that multiplies the physical spacetime volume. It is not a field and has no kinetic term and it only encodes an NG-like tension of the underlying elastic space. It is purely geometric in the sense that it couples only through $\sqrt{-g}$.
$\bullet$ A bare cosmological variable $\Lambda$ that appears as $-\Lambda\int\sqrt{-g}$.
$\bullet$ Matter with a global Weyl scaling through $\lambda$:
$S_{m}[g,\psi;\lambda]\equiv\int\sqrt{-g}\,L_{m}(\lambda^{-2}\,g_{\mu\nu},\psi)$ (3.2)
The factor $\lambda^{-2}g_{\mu\nu}$ ensures that the matter sector sees a metric that is globally Weyl related to $g_{\mu\nu}$. The variable $\lambda$ is not a field with spacetime dependence rather it is a single global variable. We will see this from the four-form sector later.
$\bullet$ Two top-form sectors,
$S_{4\text{-forms}}=\sigma\Big(\frac{\Lambda}{\mu^{4}}\Big)\int F_{4}+\hat{\sigma}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\int\hat{F}_{4}$ (3.3)
where $F_{4}=dA_{3}$ and $\hat{F}_{4}=d\hat{A}_{3}$ with $A$ and $\hat{A}$ being three form potentials (see [54, 55, 57, 58] for a working understanding of three forms). The integrals $\int F_{4}$ and $\int\hat{F}_{4}$ do not use the volume factor $\sqrt{-g}$. In essence, they are purely topological and their values depend on global fluxes and boundary conditions, not on the local metric.
c) Assumptions about the functions $\sigma$ and $\hat{\sigma}$: We assume that they are smooth and monotone enough so that the equations we derive can be inverted to view $\Lambda$ and $\lambda$ as global constants fixed by fluxes. We do not need stronger properties at this stage. We also implicitly assume boundary conditions that render the variational problem well posed for $A_{3}$ and $\hat{A}_{3}$.
d) Variation conventions: Variation with respect to $g^{\mu\nu}$ defines the total stress tensor as usual. Variation with respect to $\psi$ gives matter equations of motion that we do not need explicitly. Variation with respect to $A_{3}$ and $\hat{A}_{3}$ gives four-form equations that enforce global constraints. Variation with respect to $\Lambda$ and $\lambda$ gives global relations because the four-forms enter as Lagrange multipliers as we will demonstrate below.
In all, this is the standard sequestering setup except there is an additional NG-like term for space.
We now proceed to show that under a constant shift of the matter Lagrangian $L_{m}$ by $\Delta V$ the combination $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$ is invariant point-wise . The geometric tension $T_{s}$ is untouched by the four-form constraints because it does not couple to $\lambda$ and does not appear in the matter trace. Therefore the effective cosmological term left in the local Einstein equations equals $T_{s}$. A four-form sequestering mechanism of this type cannot neutralize the NG-like geometric tension of space.
Proof:
1.) Varying the action with respect to the three-form potentials:
Consider the term $\sigma(\Lambda/\mu^{4})\int F_{4}$, where $F_{4}=dA_{3}$. Under a variation of $A_{3}$ with compact support we have,
$\delta\left(\int\sigma\,F_{4}\right)=\int\sigma\,\delta F_{4}=\int\sigma\,d(\delta A_{3})=\int d\big(\sigma\,\delta A_{3}\big)-\int d\sigma\wedge\delta A_{3}$ (3.4)
The boundary term is $\int d(\sigma\,\delta A_{3})$. We set it to zero by appropriate boundary conditions. The remaining term is $\int d\sigma\wedge\delta A_{3}$ and given that $\delta A_{3}$ is arbitrary in the interior, the Euler–Lagrange equation is,
$d\sigma\Big(\frac{\Lambda}{\mu^{4}}\Big)=0$ (3.5)
Now because $\sigma$ depends on spacetime only through $\Lambda$, this implies that $d(\Lambda/\mu^{4})=0$ provided $\sigma^{\prime}\neq 0$. Therefore $\Lambda$ is constant in spacetime.
Repeating the same steps for $\hat{\sigma}(\lambda^{4}/\mu^{4})\int\hat{F}_{4}$, we have after varying with respect to $\hat{A}_{3}$,
$d\hat{\sigma}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)=0$ (3.6)
Assuming $\hat{\sigma}^{\prime}\neq 0$ we conclude that $d(\lambda^{4}/\mu^{4})=0$. Hence $\lambda$ is constant in spacetime and since $\lambda>0$ we can take $\lambda$ itself to be a global constant.
Therefore, the four-form sectors force $\Lambda$ and $\lambda$ to be global constants. That is, they do not propagate and they are numbers fixed by boundary data and fluxes.
2.) Varying the action with respect to the metric and identifying the local Einstein equations:
We now compute the variation of each metric dependent term with respect to $g^{\mu\nu}$ where we will be using the identity $\delta\sqrt{-g}=-\dfrac{1}{2}\sqrt{-g}\,g_{\mu\nu}\,\delta g^{\mu\nu}$. The variation of the Einstein–Hilbert part gives the usual $M_{*}^{2}G_{\mu\nu}$ on the left hand side of the field equations.
The variation of $T_{s}\int\sqrt{-g}$ gives,
$\delta\big(-T_{s}\int\sqrt{-g}\big)=-T_{s}\int\delta\sqrt{-g}=\frac{T_{s}}{2}\int\sqrt{-g}\,g_{\mu\nu}\,\delta g^{\mu\nu}$ (3.7)
This means the geometric tension contributes a stress tensor equal to $T_{s}g_{\mu\nu}$ with a minus sign when moved to the right hand side of Einstein equations.
The variation of $\Lambda\int\sqrt{-g}$ is identical in structure and it gives a negative contribution $\Lambda g_{\mu\nu}$ to the right hand side of the field equations as well.
Now we vary the matter action with respect to the metric, which reads,
$S_{m}[g,\psi;\lambda]=\int\sqrt{-g}\,\,L_{m}(\lambda^{-2}\,g,\psi)$ (3.8)
We define the physical matter stress tensor by the standard definition,
$T^{(m)}_{\mu\nu}\equiv-\frac{2}{\sqrt{-g}}\,\frac{\delta S_{m}}{\delta g^{\mu\nu}}$ (3.9)
The above definition ensures that the matter equations are consistent with diffeomorphism invariance and with the Bianchi identities. The full variation of the matter part yields,
$\delta S_{m}=-\frac{1}{2}\int\sqrt{-g}\,T^{(m)}_{\mu\nu}\,\delta g^{\mu\nu}$ (3.10)
Collecting all terms, the total variation with respect to $g^{\mu\nu}$ gives,
$\delta S=\int\sqrt{-g}\,\Big[\frac{M_{*}^{2}}{2}\,G_{\mu\nu}+\frac{1}{2}\big(T_{s}+\Lambda\big)g_{\mu\nu}-\frac{1}{2}\,T^{(m)}_{\mu\nu}\Big]\,\delta g^{\mu\nu}$ (3.11)
Stationarity for arbitrary $\delta g^{\mu\nu}$ gives,
$\frac{M_{*}^{2}}{2}\,G_{\mu\nu}+\frac{1}{2}\big(T_{s}+\Lambda\big)g_{\mu\nu}-\frac{1}{2}\,T^{(m)}_{\mu\nu}=0$ (3.12)
Multiplying by 2, the local Einstein field equations read,
$M_{*}^{2}\,G_{\mu\nu}=T^{(m)}_{\mu\nu}-\big(T_{s}+\Lambda\big)\,g_{\mu\nu}$ (E)
It is important to note that the four-form terms do not contribute to the metric equations. They are metric independent in the form we wrote, since $\int F_{4}$ and $\int\hat{F}_{4}$ do not carry $\sqrt{-g}$ or $g^{\mu\nu}$. Hence they add no local stress tensor.
So, we now have the local field equations where we see that the geometric tension $T_{s}$ and the variable $\Lambda$ both appear as volume terms and therefore enter as $(\Lambda+T_{s})\,g_{\mu\nu}$. The matter sector contributes its usual stress tensor $T^{(m)}_{\mu\nu}$.
3.) Varying the action with respect to $\Lambda$:
Only two parts of the action (3.1) depend on $\Lambda$. The volume term $\Lambda\int\sqrt{-g}$ and the top-form term $\sigma(\Lambda/\mu^{4})\int F_{4}$. Therefore, computing the variation and treating $\Lambda$ as a global number, we find,
$\delta S\big|_{\Lambda}=-\delta\Lambda\int d^{4}x\,\sqrt{-g}+\sigma^{\prime}\Big(\frac{\Lambda}{\mu^{4}}\Big)\,\frac{\delta\Lambda}{\mu^{4}}\,\int F_{4}$ (3.13)
Setting the equation above this to zero for arbitrary $\delta\Lambda$, we obtain the first global constraint,
$\int d^{4}x\,\sqrt{-g}=\frac{\sigma^{\prime}\left(\Lambda/\mu^{4}\right)}{\mu^{4}}\,\int F_{4}$ (C1)
This equation tells us that the spacetime four-volume is proportional to the four-form flux weighted by $\sigma^{\prime}$. Note that it does not fix the local metric rather it only relates global numbers.
With this, we have one global relation that involves $\Lambda$ but no local fields directly.
4.) Varying the action with respect to $\lambda$:
Again, only two parts of the action (3.1) depend on $\lambda$, which are the matter action $S_{m}$ and the top-form term $\hat{\sigma}(\lambda^{4}/\mu^{4})\int\hat{F}_{4}$. First consider the matter action which we compute $\delta S_{m}$ under a small variation of the global $\lambda$.
$\delta S_{m}\bigg|_{\lambda}=\int\sqrt{-g}\,\frac{\partial L_{m}(\lambda^{-2}\,g,\,\psi)}{\partial(\lambda^{-2}g_{\mu\nu})}\,\delta(\lambda^{-2}g_{\mu\nu})$ (3.14)
Let $h_{\mu\nu}\equiv\lambda^{-2}g_{\mu\nu}$ and its variation under a variation of $\lambda$ at fixed $g$ is,
$\delta h_{\mu\nu}=\delta(\lambda^{-2})\,g_{\mu\nu}=-2\,\lambda^{-3}\,\delta\lambda\,g_{\mu\nu}$ (3.15)
So we have,
$\int\sqrt{-g}\;\frac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\;\delta(h_{\mu\nu})=-2\,\int\sqrt{-g}\;\lambda^{-3}\,\frac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\,g_{\mu\nu}\,\delta\lambda$ (3.16)
In order to obtain $\dfrac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\,g_{\mu\nu}$ , we first vary $S_{m}$ with respect to the metric tensor $g_{\mu\nu}$, which gives,
$\delta S_{m}=-\int\sqrt{-g}\;\lambda^{-2}\,\frac{\partial L_{m}(h_{\mu\nu},\psi)}{\partial(h_{\mu\nu})}g_{\alpha\mu}\,g_{\beta\nu}\delta g^{\mu\nu}$ (3.17)
and use the definition of the matter stress energy tensor (3.9), which gives,
$T_{\mu\nu}^{(m)}=2\,\lambda^{-2}\,\frac{\partial L_{m}(h_{\mu\nu},\psi)}{\partial(h_{\mu\nu})}g_{\alpha\mu}\,g_{\beta\nu}$ (3.18)
Taking the trace, we find,
$T^{(m)}=g^{\mu\nu}\,T_{\mu\nu}^{(m)}=2\,\lambda^{-2}\,\frac{\partial L_{m}(h_{\mu\nu},\psi)}{\partial(h_{\mu\nu})}g_{\mu\nu}$ (3.19)
Solving for $\dfrac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\,g_{\mu\nu}$, we obtain,
$\dfrac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}g_{\mu\nu}=\frac{\lambda^{2}\,T^{(m)}}{2}$ (3.20)
Substituting into equation (3.16), we get,
$\int\sqrt{-g}\;\frac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\;\delta(h_{\mu\nu})=-\frac{\delta\lambda}{\lambda}\int\sqrt{-g}\,T^{(m)}$ (3.21)
Using (3.21) in (3.14), we obtain,
$\delta S_{m}\bigg|_{\lambda}=-\frac{\delta\lambda}{\lambda}\int\sqrt{-g}\,T^{(m)}$ (3.22)
Now we vary the top-form term $\hat{\sigma}(\lambda^{4}/\mu^{4})\int\hat{F}_{4}$, which gives,
$\delta\Big[\hat{\sigma}\!\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\int\hat{F}_{4}\Big]=\hat{\sigma}^{\prime}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\,\frac{4\lambda^{3}\,\delta\lambda}{\mu^{4}}\,\int\hat{F}_{4}$ (3.23)
Writing both pieces together, that is equations (3.21) and (3.23), the full variation of $S_{m}$ with respect to the global $\lambda$ then reads,
$\frac{\delta\lambda}{\lambda}\,\int\sqrt{-g}\,T^{(m)}+\hat{\sigma}^{\prime}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\,\frac{4\lambda^{3}\,\delta\lambda}{\mu^{4}}\,\int\hat{F}_{4}=0$ (3.24)
Canceling $\delta\lambda$ which is arbitrary but not zero and multiplying by $\lambda$, we obtain the second global constraint,
$\int d^{4}x\,\sqrt{-g}\,T^{(m)}=\frac{4\,\lambda^{4}}{\mu^{4}}\,\hat{\sigma}^{\prime}\left(\frac{\lambda^{4}}{\mu^{4}}\right)\,\int\hat{F}_{4}$ (C2)
We now have a second global relation that connects the spacetime integral of the matter trace to the flux of $\hat{F}_{4}$. The geometric tension $T_{s}$ does not appear because it does not depend on $\lambda$.
5.) Eliminating fluxes to relate $\Lambda$ to the spacetime average of $T^{(m)}$:
Equations (C1) and (C2) contain the fluxes $\int F_{4}$ and $\int\hat{F}_{4}$. These fluxes are constant numbers fixed by boundary conditions. We define the spacetime average of a quantity $X$ as
$\langle X\rangle=\dfrac{\int\sqrt{-g}\,X}{\int\sqrt{-g}}$ (3.25)
Combining (C1) and (C2) we may then write,
$\langle T^{(m)}\rangle=4\,\lambda^{4}\,\frac{\hat{\sigma}^{\prime}(\lambda^{4}/\mu^{4})}{\sigma^{\prime}(\Lambda/\mu^{4})}\,\frac{\int\hat{F}_{4}}{\int F_{4}}$ (3.26)
The right hand side is a pure number fixed by the values of the global variables and by the flux ratio. In the canonical sequestering normalization this relation is equivalent to,
$\Lambda=\frac{1}{4}\,\langle T^{(m)}\rangle$ (3.27)
This is the standard global identity in four-form sequestering and follows from the homogeneity in $\lambda$ together with the two Lagrange multiplier equations. Therefore, we have a global identity that expresses $\Lambda$ as one quarter of the spacetime average of the matter trace.
6.) Performing the constant shift test and identify the invariant local combination:
Consider a rigid shift of the matter Lagrangian by a constant $\Delta V$. Equivalently, we may replace $L_{m}\to L_{m}-\Delta V$ so that the added term in the action is $\int\sqrt{-g}\,\lambda^{4}\Delta V$. Using equation (3.9), we see that this simply changes the matter stress $T^{(m)}_{\mu\nu}$ and trace of matter stress $T^{(m)}$ by,
$\Delta T^{(m)}_{\mu\nu}=-\Delta V\,g_{\mu\nu},\qquad\Delta T^{(m)}=-4\,\Delta V$ (3.28)
The global relation then yields,
$\Delta\Lambda=\frac{1}{4}\,\Delta\langle T^{(m)}\rangle=-\,\Delta V$ (3.29)
Therefore the matter-vacuum local combination that appears in the Einstein equations is invariant,
$T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}\;\to\;\big(T^{(m)}_{\mu\nu}-\Delta Vg_{\mu\nu}\big)-\big(\Lambda-\Delta V\big)g_{\mu\nu}=T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$ (3.30)
All constant contributions from the matter sector cancel locally between $T^{(m)}_{\mu\nu}$ and $\Lambda g_{\mu\nu}$. This includes radiative corrections that enter as shifts of the vacuum energy in $L_{m}$.
7.) Identifying the leftover geometric NG-term:
Returning to the local Einstein equations (E),
$M_{*}^{2}\,G_{\mu\nu}=T^{(m)}_{\mu\nu}-\big(T_{s}+\Lambda\big)\,g_{\mu\nu}$ (3.31)
We rewrite the right hand side as,
$\big[T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}\big]-T_{s}\,g_{\mu\nu}$ (3.32)
The bracket is invariant under constant shifts of the matter Lagrangian. Therefore any constant vacuum energy from the matter sector is neutralized by the four-form enforced value of $\Lambda$. The term $T_{s}g_{\mu\nu}$ is untouched by this mechanism because $T_{s}$ does not depend on $\lambda$ and does not appear in the global trace equation. It is a purely geometric constant and it contributes a cosmological term to the Einstein equations that cannot be removed or adjusted by the four-form constraints. Therefore the effective cosmological term that remains in the local field equations is $T_{s}$.
Thus, the neutralization does not act on $T_{s}$. The residual cosmological constant is $T_{s}$ as claimed.
8.) General Remarks:
A. Bianchi identities and conservation: Taking the divergence of both sides of the Einstein equations, the left hand side obeys $\nabla^{\mu}G_{\mu\nu}=0$ by the contracted Bianchi identity. The right hand side has $\nabla^{\mu}T^{(m)}_{\mu\nu}=0$ because the matter action is diffeomorphism invariant and couples minimally to $g_{\mu\nu}$. The term $(T_{s}+\Lambda)\,g_{\mu\nu}$ has zero divergence because $T_{s}$ and $\Lambda$ are global constants. Hence the equations are consistent with the Bianchi identities.
B. Role of the top forms in the stress tensor: In the action (3.1) the four-form integrals carry no metric dependence and therefore do not contribute any local stress. This is why the local equations contain only $T^{(m)}_{\mu\nu}$ and the volume terms. If one used a different four-form model that included metric dependent contractions of $F_{\mu\nu\rho\sigma}$ with itself, then the four-form would carry local stress and would behave like a dynamical cosmological constant. That is the Brown–Teitelboim mechanism. Our conclusion is specific to the sequestering setup where the four-forms serve as global Lagrange multipliers rather than as local energy carriers.
C. Radiative stability: Matter loop corrections shift the matter vacuum energy by constants of arbitrarily large size. These shifts are neutralized locally because the bracket $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$ is invariant. Gravitational loops can renormalize the geometric constant $T_{s}$ since it multiplies the volume element. In an effective field theory description one treats $T_{s}$ as a renormalized parameter that must be fixed by observation. So the key point is that matter loops do not destabilize the observed value once $T_{s}$ is chosen. We will discuss this point at length in the next Section.
D. Scope of the theorem. The argument required only that the matter sector couples through a single global Weyl factor $\lambda$. It required that the geometric tension $T_{s}$ does not carry any explicit $\lambda$ dependence. If one added a non minimal coupling that made $T_{s}$ scale with $\lambda$, then the cancellation pattern would change and the result would not follow. Our setup excludes such couplings by assumption.
9.) Einstein Equations for a pure matter vacuum:
Consider a pure matter vacuum with $T^{(m)}_{\mu\nu}=-Vg_{\mu\nu}$ for some constant $V$. The trace is $T^{(m)}=-4\,V$ and the spacetime average of the trace equals $\langle T^{(m)}\rangle=-4V$. The global constraint gives $\Lambda=\dfrac{1}{4}\langle T^{(m)}\rangle=-V$. The matter-vacuum bracket therefore vanishes,
$T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}]=0$ (3.33)
and the Einstein equations reduce to,
$M_{*}^{2}G_{\mu\nu}=-T_{s}\,g_{\mu\nu}$ (3.34)
The geometric term $T_{s}$ in turn receives correction from graviton self loops and graviton-matter loops as we show in the subsequent Section. Therefore, it can be naturally identified with the residual observed Dark Energy. This conclusion is consistent with the general statement of the theorem and also makes the no geometric sequestering mechanism very transparent.
$\square$
4 Gravitational EFT, Graviton Loops and the Residual Term $T_{s}$
Our goal is to present the low energy gravitational effective action in a way that keeps the discussion compact while preserving all structural information, see [36, 37]. We use the background field method [39, 40], expand only the metric around a fixed background and keep matter fields unexpanded in this section. Background diffeomorphism invariance then forces the one particle irreducible functional $\Gamma$ to be a sum of background covariant operators. This locks in the tensor structure of counterterms and of the finite Wilsonian terms. Sequestering relation will then isolate the surviving constant source as the geometric volume tension $T_{s}$.
We introduce a background metric $\bar{g}_{\mu\nu}$ and the fluctuation
$g_{\mu\nu}\equiv\bar{g}_{\mu\nu}+\kappa\,h_{\mu\nu},\qquad\kappa^{2}\equiv 32\pi G\equiv\frac{4}{M_{*}^{2}}$ (4.1)
The action is still given by (3.1),
$S[g,\psi;\Lambda,\lambda]=\int d^{4}x\,\sqrt{-g}\,\Big(\tfrac{M_{*}^{2}}{2}\,R-T_{s}-\Lambda\Big)+S_{m}[\lambda^{-2}g,\psi]+S_{\text{4-forms}}[\Lambda,\lambda]$ (4.2)
where the four form sector enforces the global constant constrains on $\Lambda$ and $\lambda$.
We couple a source $J^{\mu\nu}$ to the metric fluctuation and define the gauge fixed partition function,
$Z[\bar{g},J]=\int\mathcal{D}h\,\mathcal{D}c\,\mathcal{D}\bar{c}\,\mathcal{D}\psi\;e^{i\Big(S[\bar{g}+\kappa h,\psi]+S_{\text{gf}}[\bar{g},h]+S_{\text{gh}}[\bar{g},h,c,\bar{c}]+\int d^{4}x\,\sqrt{-\bar{g}}\,J^{\mu\nu}h_{\mu\nu}\Big)}$ (4.3)
where $S_{\text{gf}}[\bar{g},h]$ and $S_{\text{gh}}[\bar{g},h,c,\bar{c}]$ are the gravitational gauge fixing and ghost Lagrangians respectively as discussed below.
We now define the connected generator,
$W[\bar{g},J]\equiv-\,i\ln Z[\bar{g},J]$ (4.4)
The mean fluctuation field is therefore,
$\hat{h}_{\mu\nu}(x)\equiv\frac{1}{\sqrt{-\bar{g}}}\frac{\delta W[\bar{g},J]}{\delta J^{\mu\nu}(x)}$ (4.5)
We have introduced $\hat{h}$ in order to Legendre transform to the one-particle-irreducible functional. This produces vertex functions, clean Ward identities and the tadpole condition that fixes the background. The Legendre transform gives,
$\Gamma[\bar{g},\hat{h}]\equiv W[\bar{g},J]-\int d^{4}x\,\sqrt{-\bar{g}}\,J^{\mu\nu}\hat{h}_{\mu\nu},\qquad\frac{1}{\sqrt{-\bar{g}}}\frac{\delta\Gamma}{\delta\hat{h}_{\mu\nu}}=-J^{\mu\nu}$ (4.6)
Setting $J=0$ enforces the stationarity condition
$\left.\frac{\delta\Gamma[\bar{g},\hat{h}]}{\delta\hat{h}_{\mu\nu}}\right|_{J=0}=0$ (4.7)
By construction we choose the background so that $\hat{h}=0$ solves (4.7). The vacuum functional is then,
$\Gamma[\bar{g}]\equiv\Gamma[\bar{g},\hat{h}]\big|_{\hat{h}=0}$ (4.8)
This object generates the quantum-corrected field equations for the background metric.
We now fix diffeomorphisms with a background covariant gauge. A standard choice is de Donder gauge built from $\bar{\nabla}$, where $\bar{\nabla}$ (the covariant derivative defined on background metric $\bar{g}$) is the unique torsion–free, metric–compatible connection satisfying,
$\bar{\nabla}_{\rho}\bar{g}_{\mu\nu}=0,\qquad\bar{T}^{\rho}{}_{\mu\nu}\equiv\bar{\Gamma}^{\rho}{}_{\mu\nu}-\bar{\Gamma}^{\rho}{}_{\nu\mu}=0$ (4.9)
and where,
$\bar{\Gamma}^{\rho}{}_{\mu\nu}=\tfrac{1}{2}\,\bar{g}^{\rho\sigma}\big(\partial_{\mu}\bar{g}_{\nu\sigma}+\partial_{\nu}\bar{g}_{\mu\sigma}-\partial_{\sigma}\bar{g}_{\mu\nu}\big)$ (4.10)
Using de Donder’s gauge, we have,
$F_{\mu}^{\text{grav}}=\bar{\nabla}^{\nu}h_{\mu\nu}-\frac{1}{2}\bar{\nabla}_{\mu}h,\qquad\mathcal{L}_{\text{gf}}^{\text{grav}}=-\frac{1}{2\alpha_{g}}\,\sqrt{-\bar{g}}\,\bar{g}^{\mu\nu}F_{\mu}F_{\nu}$ (4.11)
The Faddeev–Popov ghost action is then [41],
$\Delta_{\ \nu}^{\mu}=\bar{\nabla}^{2}\delta^{\mu}_{\ \nu}+\bar{R}^{\mu}_{\ \nu},\qquad\mathcal{L}_{\text{gh}}^{\text{grav}}=-\sqrt{-\bar{g}}\,\bar{c}_{\mu}\,\Delta_{\ \nu}^{\mu}\,c^{\nu}$ (4.12)
Let $\xi^{\mu}(x)$ generate an infinitesimal diffeomorphism. Background diffeomorphisms act as ordinary diffeomorphisms on $\bar{g}$ and as tensorial Lie derivatives on all quantum fields and on the source $J$,
$\delta_{\text{bg}}\bar{g}_{\mu\nu}=\bar{\nabla}_{\mu}\xi_{\nu}+\bar{\nabla}_{\nu}\xi_{\mu},\qquad\delta_{\text{bg}}h_{\mu\nu}=\mathcal{L}_{\xi}h_{\mu\nu},$
$\delta_{\text{bg}}c^{\mu}=\mathcal{L}_{\xi}c^{\mu},\qquad\delta_{\text{bg}}\bar{c}_{\mu}=\mathcal{L}_{\xi}\bar{c}_{\mu},\qquad\delta_{\text{bg}}J^{\mu\nu}=\mathcal{L}_{\xi}J^{\mu\nu}$ (4.13)
Because $S_{\text{gf}}$ and $S_{\text{gh}}$ are built from $\bar{g}$ and $\bar{\nabla}$, they are invariant under $\delta_{\text{bg}}$. The classical action $S[\bar{g}+\kappa h,\psi]$ is diffeomorphism invariant by construction. Furthermore, we adopt a background-covariant regularization and measure, for instance dimensional regularization [38], so that $Z$ and $W$ are invariant when $J$ transforms as a background tensor. The Legendre transform then gives the Ward identity,
$\delta_{\text{bg}}\Gamma[\bar{g},\hat{h}]=0,\qquad\delta_{\text{bg}}\hat{h}_{\mu\nu}=\mathcal{L}_{\xi}\hat{h}_{\mu\nu}$ (4.14)
If we set $\hat{h}=0$, we obtain $\delta\,\Gamma[\bar{g}]=0$, which implies that $\Gamma[\bar{g}]$ is invariant under diffeomorphisms of $\bar{g}$ and therefore it can only be an integral of scalar densities built from $\bar{g}_{\mu\nu}$, the Riemann tensor of $\bar{g}$ and background covariant derivatives. This is the structural reason why the effective action organizes into background-covariant operators.
Equation (4.14) is the reason why the effective action admits an expansion in background covariant operators. Every local counterterm and every local Wilsonian term (see [51, 42, 43, 44]) is an integral of $\sqrt{-\bar{g}}$ times a scalar built from $\bar{R}$, $\bar{R}_{\mu\nu}$, $\bar{R}_{\mu\nu\rho\sigma}$, and $\bar{\nabla}$ and it takes the universal form for the local part,
$\Gamma^{\text{local}}[\bar{g}]=\int d^{4}x\,\sqrt{-\bar{g}}\,\Big[c_{0}(\mu)+\frac{1}{2}\,c_{2}(\mu)\,R+c_{4}^{(1)}(\mu)\,R^{2}+c_{4}^{(2)}(\mu)\,R_{\mu\nu}\,R^{\mu\nu}+\cdots\Big]$ (4.15)
The non-local part contains structures such as $R\log(-\bar{\Box}/\mu^{2})R$ [49, 45, 50]. These modify long distance kernels but do not act as constant background sources such as dark energy because they contain derivatives.
In particular, the coefficient $c_{0}(\mu)$ renormalizes the volume term, that is $T_{s}$, the coefficient $c_{2}(\mu)$ renormalizes the Einstein–Hilbert coupling (Newton’s constant $G$), the $c_{4}^{(i)}(\mu)$ term renormalizes curvature squared operators and higher terms contain more derivatives and are suppressed at low curvature.
The parameters $c_{n}(\mu)$ run with the renormalization scale $\mu$. We write,
$\mu\,\frac{d}{d\mu}\,c_{0}(\mu)=\beta_{0}(\mu),\qquad\mu\,\frac{d}{d\mu}\,c_{2}(\mu)=\beta_{2}(\mu),\qquad\mu\,\frac{d}{d\mu}\,c^{(i)}_{4}(\mu)=\beta^{(i)}_{4}(\mu),\quad\text{etc.}$ (4.16)
The precise values depend on the field content and on the gauge fixing and on the regulator.
We now combine tree level and loop corrected coefficients that multiply the same operators. We define renormalized parameters at the scale $\mu$ by
$T_{s}(\mu)$ $\equiv T_{s}^{\text{bare}}-c_{0}(\mu),$
$M_{*}^{2}(\mu)$ $\equiv M_{*\,\text{bare}}^{2}+c_{2}(\mu),$
$\alpha_{1,2}(\mu)$ $\equiv c_{4}^{(1,2)}(\mu)$ (4.17)
The local gravitational part of the effective action can therefore be written as,
$\Gamma^{\text{grav}}_{\text{local}}=\int d^{4}x\,\sqrt{-g}\,\Big(-T_{s}(\mu)+\frac{M_{*}^{2}(\mu)}{2}R+\alpha_{1}(\mu)R^{2}+\alpha_{2}(\mu)R_{\mu\nu}R^{\mu\nu}+\cdots\Big)$ (4.18)
Varying and adding the matter part gives the local field equations,
$M_{*}^{2}(\mu)\,G_{\mu\nu}=[T^{(m)}_{\mu\nu}-\Lambda\,g_{\mu\nu}]-T_{s}(\mu)\,g_{\mu\nu}+\mathcal{H}_{\mu\nu}[R^{2},\dots]$ (4.19)
where $\mathcal{H}_{\mu\nu}$ collects higher derivative tensors from curvature squared terms. These tensors are suppressed at low curvature and do not act like a constant source. The combination $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$ is invariant under rigid shifts $L_{m}\to L_{m}-\Delta V$ in the sequestering setup, hence constant vacuum energy inside $S_{m}$ cancels out of the local equations.
So in the field equation (4.19), the net constant vacuum source after sequestering is
$\rho_{\rm vac}(\mu)\,g_{\mu\nu}=-T_{s}(\mu)\,g_{\mu\nu}$ (4.20)
At late times one can match this constant source to the observed dark energy density,
$\rho_{\Lambda}^{\rm obs}=T_{s}(\mu_{\rm IR})$ (4.21)
where $\mu_{\rm IR}$ is an infrared matching scale appropriate for cosmology. This identification fixes one linear combination of low energy parameters using data.
5 Discussion
We developed a minimal framework in which space is modeled as an elastic three membrane that fills spacetime. The dynamics are captured by an intrinsic Nambu–Goto volume term $S_{s}=-T_{s}\int\sqrt{-g}\,d^{4}x$. This fixes the geometric stress tensor to $T^{(s)}_{\mu\nu}=-T_{s}\,g_{\mu\nu}$ and identifies the membrane tension with a uniform vacuum energy of space. We then embedded this term into a matter vacuum sequester framework that employs global four form constraints to cancel constant contributions from the matter sector. The variation shows that the global variable adjusts to neutralize the matter vacuum energy while a purely geometric unit operator (the operator which behaves like vacuum energy) carries no dependence on that variable and is not in the cancellation channel. As a result the local field equations retain a residual contribution $T_{s}\,g_{\mu\nu}$. We interpret this as the effective cosmological term.
Subsequently we performed the path integral over the gravitational fluctuation $h_{\mu\nu}$ and over the associated Faddeev–Popov ghosts only. We did not integrate over matter fluctuations so as to not “recompute” the pure matter loops. Their effect is assumed to have already been absorbed into the renormalized matter functional $S_{m}[\lambda^{-2}g,\psi]$ at the matching scale $\mu$.
The outcome was a background diffeomorphism invariant one particle irreducible action $\Gamma[\bar{g}]$ with the structure shown in equation (4.15). The local part is an infinite covariant tower built from the curvature and its derivatives. The non-local part contains form factors that never reduce to a pure volume term and so do not influence vacuum energy. The geometric volume tension $T_{s}(\mu)$ arises from the renormalized coefficient of the unit operator and enters the field equations. This separation makes clear which ingredients can shift $T_{s}(\mu)$ and which ingredients only change higher derivative responses of the geometry.
If one wishes to include matter-graviton mixed loops explicitly, one can extend the path integral to the matter fields fluctuations as well. The background field method and general covariance then enforce the same covariant operator basis that appears in equation (4.15). Mixed matter-graviton diagrams that contain at least one internal graviton line are local in the ultraviolet. They therefore renormalize the same coefficients that multiply curvature invariants as in equation (4.15). No new local scalar of lower canonical mass dimension exists beyond the unit operator and the Ricci scalar term. As a result the only change is in the numerical values of the Wilson coefficients.
Concretely, mixed matter–graviton loops shift the same $c_{0}(\mu)$ and therefore in turn shift $T_{s}(\mu)\equiv T_{s}^{\rm bare}-c_{0}(\mu)$. They also shift $c_{2}(\mu)$ which renormalizes the constant $G$ along with shifting the coefficients of curvature squared invariants and beyond. The tensorial form of each contribution is fixed by background diffeomorphisms and does not depend on the details of the gauge fixing. It can further be arranged so that pure matter functional $S_{m}[\lambda^{-2}g,\psi]$ will still continue to encode matter self interactions and pure matter loops at the chosen matching scale. The combined result is a single covariant effective action with renormalized coefficients $c_{0}(\mu),c_{1}(\mu),\cdots$ that capture both gravitational and mixed contributions in a uniform way. Therefore, the generalization to mixed matter-graviton graphs is straightfoward.
With this in mind, one can then conveniently regard $T_{s}(\mu)$ as the fully renormalized geometric tension that already includes contributions from pure graviton self loops and from mixed matter–graviton loops because as mentioned before, extending the path integral from metric fluctuations to metric plus matter fluctuations only shifts the Wilson coefficients in the same covariant operator basis. The definition of $T_{s}(\mu)$ as the renormalized coefficient of the unit operator therefore captures all such shifts in a single parameter.
One might try to redefine $\Lambda^{\prime}\equiv\Lambda-T_{s}$ so that the sequestering constraint cancels both the matter vacuum energy and the geometric volume term. This identification is physically not possible because in the standard sequestering framework which we are considering the global variables and four-forms are wired so that the variational constraint removes constant shifts of the matter Lagrangian, while the purely gravitational sector is inert under that constraint. An independent geometric unit operator $T_{s}\int\sqrt{-g}$ is therefore not neutralized unless one explicitly extends the mechanism to couple the same global sector to that operator. To also cancel $T_{s}$, one would need a different theory, one which couples a new rigid variable to a geometric density. This represents a different sequestering scheme and whereas we are working in the minimal framework which requires just two global Lagrange multipliers $\lambda$ and $\Lambda$. One can also see why such an identification is not possible in the minimal case because the operator $T_{s}\int\sqrt{-g}$ is blind to $\lambda$ and therefore cannot contribute to the $\lambda$ constrain.
So it is correct to keep $\Lambda$ as the variable that enforces the global matter constraint and we keep $T_{s}(\mu)$ as an intrinsic geometric coupling that lives in the gravitational effective action. In the minimal sequester framework, the added advantage of having an un-sequestered geometric term is that after fixing covariant operator basis using the background field method, the term $T_{s}(\mu)$ automatically emerges as the renormalized coefficient of the unit operator after graviton and mixed matter–graviton loops are included. Nonlocal form factors appear but they do not produce a pure volume term and therefore do not modify this identification. The constant source in the field equations can therefore naturally be identified as $\rho_{\rm cav}(\mu)=T_{s}(\mu)$. At an infrared scale $\mu_{\rm IR}$ appropriate for cosmology one can then match $\rho_{\rm vac}(\mu_{\rm IR})$ to the observed dark energy density. This retains the virtue of sequestering for matter vacuum energy while leaving a natural geometric volume term that can account for the residual dark energy. It also avoids promoting geometric couplings to global multipliers and keeps conservation and locality manifest through the Bianchi identity.
6 Conclusion
We have modeled space as an elastic three membrane and identified its uniform volume tension with a geometric unit operator that contributes a stress proportional to the metric. The Nambu–Goto term $S_{s}=-T_{s}\int\sqrt{-g}\,d^{4}x$ fixes $T^{(s)}_{\mu\nu}=-T_{s}g_{\mu\nu}$ and allows a direct comparison with vacuum energy. Embedding this structure into a four–form sequester framework, we showed that global constraints neutralize constant matter contributions inside the invariant combination $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$, while the geometric unit operator does not couple to the global variable and therefore cannot be removed. The local field equations then take the renormalized effective form,
$M_{*}^{2}(\mu)\,G_{\mu\nu}=\bigl[T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}\bigr]-T_{s}(\mu)\,g_{\mu\nu}+H_{\mu\nu}[R^{2},\nabla R,\ldots]$
with higher–derivative tensors suppressed at low curvature. Using the background field method we argued that graviton loops and mixed matter–graviton loops renormalize the unit operator, so $T_{s}(\mu)$ runs with the scale. The residual constant source may be identified with dark energy at an infrared matching scale relevant for cosmology.
An immediate next step is to extend the elastic space picture by introducing bulk modulus and bulk viscosity to space. In a homogeneous expansion a nonzero bulk viscosity produces an effective negative pressure that can evolve with the Hubble rate. This offers a natural route to time dependent dark energy and can help account for DESI indications of evolving dark energy. We will develop this viscous elastic space framework next.
7 Acknowledgment
The author acknowledges the use of ChatGPT-5 (August 2025 version) to edit and refine tone and clarity in selected sections of this manuscript. The model was used only to polish and rephrase text supplied by the author. All LLM assisted passages were reviewed and edited by the author, exercising the final editorial judgment.

### v3

@@
-The Tension of Space as Dark Energy: A No Geometric Sequestering Theorem for Minimal Sequesters
+The Tension of Space as Dark Energy: Dynamics and Phenomenology
 Abstract
-We model space as an elastic membrane and identify its uniform tension $T_{s}$ with the vacuum energy density of space.
-The central result is a no geometric sequestering theorem.
-In any minimal matter vacuum sequester framework where a global constraint cancels constant contributions from the matter sector by adjusting a global variable, a purely geometric unit operator remains intact.
-Concretely, a volume term of the form $S_{s}=-T_{s}\int\sqrt{-g}\,d^{4}x$ does not couple to the global variable that enforces the matter cancellation, and it survives unchanged in the local field equations.
-We establish this result within a covariant effective field theory using the background field method.
-The analysis shows that the coefficient $T_{s}$ is renormalized by graviton loops and in general by mixed matter–graviton loops and therefore runs with the renormalization scale $\mu$.
-The running defines an effective cosmological term $T_{s}(\mu)$ that is present even after the matter vacuum energy has been neutralized by the sequester constraint.
-In particular, we find that $T_{s}(\mu)$ sources $G_{\mu\nu}=-T_{s}(\mu)\,g_{\mu\nu}+\cdots$.
-The tension term $T_{s}(\mu)$ can then be understood as residual dark energy in the minimal sequestering framework.
+We propose a phenomenological framework in which the observed late-time dark-energy sector is interpreted as the intrinsic tension of space itself.
+Motivated by recent observational hints that the dark-energy equation of state may exhibit mild low-redshift evolution, we begin from an intrinsic membrane description of spacetime and show that a uniform space tension contributes to the gravitational field equations with precisely the tensor structure of vacuum energy.
+We then consider its Dirac–Born–Infeld completion, which naturally introduces a hidden Abelian $U(1)$ gauge sector on the membrane of space.
+Within this setting, we study a late-time hidden symmetry-breaking transition $U(1)_{h}\to\mathbb{Z}_{n}$, which reorganizes hidden magnetic energy into a coarse-grained flux-tube reservoir.
+If this reservoir exchanges energy with the dynamical part of the space-tension sector, the effective dark-energy density becomes time dependent and acquires a nontrivial equation of state.
+At the background level, the model yields a simple proof-of-concept realization of mild running dark energy and admits a transient crossing of the phantom divide.
+We compare the resulting evolution with a compressed observational benchmark in the effective Chevallier–Polarski–Linder plane.
+Although the minimal single-reservoir realization does not exactly reproduce the benchmark target, it generates a phenomenologically relevant low-redshift evolution and points toward a late hidden-sector transition together with a substantial hidden defect population.
+The construction should therefore be viewed not as a complete theory of dark energy, but as a physically transparent proof of concept.
 1 Introduction
@@
 and found them to be dimmer than expected in a decelerating model.
-The inferred expansion is accelerating.
+The conclusion was that the expansion of space is accelerating.
 Independent evidence from the Cosmic Microwave Background anisotropies by DASI [8], WMAP [9], and Planck [10], and from Baryon Acoustic Oscillations [11, 12], supports this conclusion.
@@
 Dark Energy and the Cosmological Constant Problem
-The vacuum contribution that drives late time acceleration is often called dark energy.
+This vacuum contribution that drives late time acceleration is often called dark energy.
 It is dark because it does not radiate or absorb in a way we can directly detect, and its evidence comes from dynamics such as the acceleration rate.
@@
 Many ideas have been explored.
-Representative proposals include quintessence [15, 16], k-essence [17, 18], modified gravity [19, 20, 21, 22, 23], the Dvali Gabadadze Porrati braneworld [24, 25], Chaplygin gas [26], holographic dark energy models [27, 28, 29], quantum cosmology [54], dynamical relaxation of the Cosmological Constant via membrane nucleation events of top forms [57] and multiple top forms [58].
+Representative proposals include quintessence [15, 16], k-essence [17, 18], modified gravity [19, 20, 21, 22, 23], the Dvali Gabadadze Porrati braneworld [24, 25], Chaplygin gas [26], holographic dark energy models [27, 28, 29], quantum cosmology [30, 31, 32], dynamical relaxation of the Cosmological Constant via membrane nucleation events of top forms [33] and multiple top forms [34].
 For broad reviews see [35, 36].
-Our goal here is to give a simple and intuitive origin for the cosmological term, and hence for dark energy, using only basic ideas that are compatible with but not restricted to string theory.
-The key input is the physics of extended objects with tension.
-We use this as a guide to show that a stress which is proportional to the metric arises naturally when a three dimensional medium with a constant volume tension fills space.
-We then argue that ordinary space itself can be modeled as such a medium.
-The resulting stress tensor has the same form as vacuum energy.
-We begin with two parallel statements. For an infinitely extended fundamental three brane one finds,
-$T_{\mu\nu(3)}=-T_{3}\,g_{\mu\nu}$ (1.10)
-where $T_{3}$ is the brane tension with dimensions of energy density. If instead we treat space itself as an elastic three brane with an intrinsic volume tension $T_{s}$, then,
-$T_{\mu\nu(s)}=-T_{s}\,g_{\mu\nu}$ (1.11)
-Both expressions have the same structure as the vacuum stress tensor discussed earlier.
-The second statement is stronger conceptually because it identifies space itself as the source of the vacuum energy.
-We will analyze how a geometric tension of space and a four form sequestering structure interact with the vacuum sector and we will show that while matter vacuum energy can be neutralized by global constraints, a purely geometric NG-like tension cannot be removed.
-The residual cosmological term is then a geometric constant that we denote by $T_{s}$ and we identify it with Dark Energy.
+[…]
+If this observational trend persists, then the late-time cosmic acceleration may not be sourced by a perfectly rigid cosmological constant, but rather by an effective vacuum sector that evolves slowly over cosmological time.
+This motivates the search for physically transparent phenomenological descriptions of dynamical dark energy.
+In the present work we explore one such possibility.
+Instead of introducing an ad hoc dark-energy fluid from the outset, we interpret the observed vacuum sector as the intrinsic tension of space itself.
+The purpose of the present work is more modest than solving the full cosmological constant problem.
+We do not attempt to explain from first principles why the absolute vacuum scale is so small.
+We assume that such a suppression may arise through mechanisms such as vacuum-energy sequestering, Hawking-type Euclidean arguments for the cosmological term, four-form flux discretuum scenarios in which many flux sectors allow an exceptionally small effective $\Lambda$ to emerge, or treating the vacuum as a self-sustained superfluid-like medium that can dynamically adjust its effective vacuum energy toward equilibrium, see for example [30, 31, 32, 42, 43, 44, 45, 46, 47, 48, 49, 51, 50, 52, 53, 54].
+Instead, we investigate whether the late-time dark-energy sector that survives after all ultraviolet questions are set aside may admit a physically transparent effective description in terms of a running tension of space.
+This viewpoint allows one to connect the vacuum-like contribution in Einstein’s equations to an intrinsic geometric property of spacetime, and then to ask how additional hidden-sector structure might perturb that contribution away from exact constancy which may result in a dynamical equation of state for dark energy.
+The roadmap of the paper is as follows.
+In Section 2 we formulate the intrinsic membrane picture in its minimal form.
+Starting from the standard Nambu–Goto action for extended objects, we specialize to the case in which the three-brane is identified with physical space itself, and show explicitly that a uniform space tension contributes to the field equations with exactly the tensor structure of vacuum energy.
+We then generalize this intrinsic description to its Dirac–Born–Infeld completion, which naturally introduces a hidden Abelian gauge sector on the membrane of space.
+In Section 3 we investigate a late-time effective origin for the running of the space tension.
+We show how a hidden $U(1)_{h}\to\mathbb{Z}_{n}$ breaking can generate a coarse-grained flux-tube reservoir, how the total tension sector may be split into a rigid background piece and a dynamical contribution, and how energy exchange between the two leads to a time-dependent effective dark-energy equation of state.
+In Section 4 we compare the resulting running-tension evolution with the effective CPL benchmark suggested by current late-time cosmological data, using a compressed Gaussian description of the $(w_{0},w_{a})$ posterior.
+Finally, in Section 5 we discuss the physical interpretation of the construction, its limitations, and possible directions for future work.
+A complementary companion work, Ref.
+[55], develops an alternative late-time realization of the same general picture of dynamical dark energy, in which the residual tension of space is supplemented not by a hidden defect reservoir but by a viscous longitudinal phonon fluid living on the elastic brane.
+We now begin with the minimal intrinsic description in which space is treated as an elastic membrane with a Nambu–Goto type volume action.
 2 Modeling Space as an Elastic Membrane
 2.1 Review of Nambu–Goto Dynamics for Strings and Branes
-A fundamental string in $1+1$ dimensions has worldsheet action,
+A fundamental string in $1+1$ dimensions has worldsheet action [56],
 $S_{1}=-T_{1}\int d^{2}\xi\,\sqrt{-\det(\gamma_{\alpha\beta})}$ (2.1)
@@
 $\gamma_{\alpha\beta}=\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}g_{\mu\nu}$ (2.2)
-For a $d$ brane in (d+1)(d{+}1) dimensions the natural generalization is [32, 33, 34],
+For a $d$ brane in (d+1)(d{+}1) dimensions the natural generalization is [57, 58],
 $S_{d}=-T_{d}\int d^{d+1}\xi\,\sqrt{(-1)^{d}\det(\gamma_{\alpha\beta})}$ (2.3)
@@
 $S_{3}=-T_{3}\int d^{4}\xi\,\sqrt{-\det(\gamma_{\alpha\beta})}$ (2.4)
-2.2 Stress Tensor of an Infinitely Extended Three Brane
+2.2 Stress Tensor of an Infinitely Extended $D_{3}$ Brane
 The stress tensor that couples to the ambient metric is defined by,
@@
 $T_{\mu\nu(3)}=-T_{3}\ g_{\mu\nu}$ (2.19)
-as claimed in Sec.
-1.
-This matches the expected units because $T_{3}$ is an energy density.
 Comparing with equation (1.5) in the Introduction we see that a uniform three brane tension $T_{3}$ behaves gravitationally like a vacuum energy density $\rho_{\mathrm{vac}}c^{2}$.
-2.3 Space as the Three Brane
+2.3 Space as a $D_{3}$ Brane
 The previous derivation invites a sharper hypothesis.
@@
 Hence, dimensionally $\rho_{\mathrm{vac}}c^{2}]=[T_{s}]=[T_{3}]$, and $S_{s}$ given by equation (2.20) exactly mimics $S_{\Lambda}$ at the level of local dynamics.
-We now look at the stress tensor that follows from the elastic action of space equation (2.20). On varying $S_{s}$, we obtain,
+We now look at the stress tensor that follows from the elastic action of space, that is equation (2.20). On varying $S_{s}$, we obtain,
 $\delta S_{s}=-T_{s}\int d^{4}x\,\delta\sqrt{-g}=-\frac{T_{s}}{2}\int d^{4}x\,\sqrt{-g}\,g^{\mu\nu}\delta g_{\mu\nu}$ (2.24)
@@
 $T_{\mu\nu(s)}=-T_{s}\,g_{\mu\nu}$ (2.28)
-Thus the elastic model of space reproduces the vacuum stress tensor with the identification $T_{s}=\rho_{\mathrm{vac}}c^{2}$.
-Therefore, the Nambu–Goto structure ensures that a uniform volume tension produces a stress proportional to the metric.
-For an infinitely extended three brane this gives $T_{\mu\nu(3)}=T_{3}g_{\mu\nu}$.
-When we elevate space itself to the role of the three brane with intrinsic volume tension $T_{s}$, the same computation yields $T_{\mu\nu(s)}=-T_{s}g_{\mu\nu}$.
-Hence the cosmological term can be interpreted as the elastic volume tension of space.
-In the next section we will couple this picture to a four form sequestering framework and show that while matter vacuum energy can be neutralized by global constraints, a geometric volume tension cannot be removed.
-The residual dark energy is then precisely this $T_{s}$.
-It is important to mention here that we do not assume that physical space is embedded in a higher dimensional ambient manifold.
+Thus the elastic model of space reproduces the vacuum stress tensor with the identification $T_{s}=\rho_{\mathrm{vac}}\,c^{2}$. Equivalently, we have,
+$T_{s}=\frac{\Lambda c^{4}}{8\pi G}\,.$ (2.29)
+In the present work, we take the next step and identify this intrinsic tension of space directly with the observed late-time vacuum sector.
+More precisely, $T_{s}$ is not interpreted as the naive bare zero-point energy of all ultraviolet degrees of freedom.
+Rather, it is understood as the small residual vacuum energy density that remains after the ultraviolet physics responsible for suppressing the enormous microscopic contributions has already done its job as mentioned earlier.
+Our standpoint here is deliberately agnostic about which ultraviolet mechanism is ultimately realized in nature.
+We simply take the observed small cosmological term as an infrared input and identify it with the effective volume tension of space itself.
+In this sense, $T_{s}$ should be viewed as the residual vacuum sector that survives after the ultraviolet vacuum-energy problem has been dealt with, rather than as a first-principles derivation of that suppression within the present model.
+Therefore, for $\Lambda\approx 1.09\times 10^{-52}\,\mathrm{m}^{-2}$ [59], equation (2.29) gives the tension of space as,
+$T_{s}\approx 5.249\times 10^{-10}\;\mathrm{J/m^{3}}$ (2.30)
+Furthermore, it is important to mention here that we do not assume that physical space is embedded in a higher dimensional ambient manifold.
 The elastic membrane language in this paper is an intrinsic description and all observables are built from the intrinsic metric and its derived tensors.
@@
 For these reasons we avoid committing to an embedding and keep the framework intrinsic from the outset.
-3 A No-Geometric-Neutralization Theorem for Sequestering Four–Forms
-We now work in natural units $\hbar=c=1$ in $3+1$ dimensions and introduce the sequester action which includes the usual Einstein–Hilbert term, a fundamental $3$-form Nambu-Goto action for space (essentially treating it as an elastic membrane), a matter sector that couples to a globally scaled metric through a single global variable $\lambda$ and two four-form sectors $F_{4}$ and $\hat{F}_{4}$ that enforce global constraints, see [61, 62, 63, 64, 65] for the detailed framework.
-The action has the following form,
-$S=$ $\int d^{4}x\,\sqrt{-g}\,\Big(\frac{M_{*}^{2}}{2}\,R-T_{s}-\Lambda\Big)+\int d^{4}x\,\sqrt{-g}\,L_{m}(\lambda^{-2}\,g_{\mu\nu},\psi)$
-$+\sigma\Big(\frac{\Lambda}{\mu^{4}}\Big)\,\int F_{4}+\hat{\sigma}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\,\int\hat{F}_{4}$ (3.1)
-with $M_{*}^{2}=1/8\pi G$.
-Note that we are working in a convention where the units of $\Lambda$ are $\Lambda]=\rm mass^{4}$ (energy density).
-The matter fields are collected into $\psi$ and the scale $\mu$ is a fixed reference scale.
-A note on notation, content of the action and minimal assumptions are in order,
-a) We work with a Lorentzian metric $g_{\mu\nu}$ of signature (+++)(-+++). The Levi–Civita connection is metric compatible and curvature tensors have the usual definitions.
-b) The total action includes the following pieces:
-$\bullet$ Gravity: $\int\sqrt{-g}\,\frac{M_{*}^{2}}{2}\,R$.
-$\bullet$ Geometric tension of space: $-\,T_{s}\int\sqrt{-g}$.
-One may think of $T_{s}$ as a constant that multiplies the physical spacetime volume.
-It is not a field and has no kinetic term and it only encodes an NG-like tension of the underlying elastic space.
-It is purely geometric in the sense that it couples only through $\sqrt{-g}$.
-$\bullet$ A bare cosmological variable $\Lambda$ that appears as $-\Lambda\int\sqrt{-g}$.
-$\bullet$ Matter with a global Weyl scaling through $\lambda$:
-$S_{m}[g,\psi;\lambda]\equiv\int\sqrt{-g}\,L_{m}(\lambda^{-2}\,g_{\mu\nu},\psi)$ (3.2)
-The factor $\lambda^{-2}g_{\mu\nu}$ ensures that the matter sector sees a metric that is globally Weyl related to $g_{\mu\nu}$.
-The variable $\lambda$ is not a field with spacetime dependence rather it is a single global variable.
-We will see this from the four-form sector later.
-$\bullet$ Two top-form sectors,
-$S_{4\text{-forms}}=\sigma\Big(\frac{\Lambda}{\mu^{4}}\Big)\int F_{4}+\hat{\sigma}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\int\hat{F}_{4}$ (3.3)
-where $F_{4}=dA_{3}$ and $\hat{F}_{4}=d\hat{A}_{3}$ with $A$ and $\hat{A}$ being three form potentials (see [54, 55, 57, 58] for a working understanding of three forms).
-The integrals $\int F_{4}$ and $\int\hat{F}_{4}$ do not use the volume factor $\sqrt{-g}$.
-In essence, they are purely topological and their values depend on global fluxes and boundary conditions, not on the local metric.
-c) Assumptions about the functions $\sigma$ and $\hat{\sigma}$: We assume that they are smooth and monotone enough so that the equations we derive can be inverted to view $\Lambda$ and $\lambda$ as global constants fixed by fluxes.
-We do not need stronger properties at this stage.
-We also implicitly assume boundary conditions that render the variational problem well posed for $A_{3}$ and $\hat{A}_{3}$.
-d) Variation conventions: Variation with respect to $g^{\mu\nu}$ defines the total stress tensor as usual.
-Variation with respect to $\psi$ gives matter equations of motion that we do not need explicitly.
-Variation with respect to $A_{3}$ and $\hat{A}_{3}$ gives four-form equations that enforce global constraints.
-Variation with respect to $\Lambda$ and $\lambda$ gives global relations because the four-forms enter as Lagrange multipliers as we will demonstrate below.
-In all, this is the standard sequestering setup except there is an additional NG-like term for space.
-We now proceed to show that under a constant shift of the matter Lagrangian $L_{m}$ by $\Delta V$ the combination $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$ is invariant point-wise .
-The geometric tension $T_{s}$ is untouched by the four-form constraints because it does not couple to $\lambda$ and does not appear in the matter trace.
-Therefore the effective cosmological term left in the local Einstein equations equals $T_{s}$.
-A four-form sequestering mechanism of this type cannot neutralize the NG-like geometric tension of space.
-Proof:
-1.) Varying the action with respect to the three-form potentials:
-Consider the term $\sigma(\Lambda/\mu^{4})\int F_{4}$, where $F_{4}=dA_{3}$. Under a variation of $A_{3}$ with compact support we have,
-$\delta\left(\int\sigma\,F_{4}\right)=\int\sigma\,\delta F_{4}=\int\sigma\,d(\delta A_{3})=\int d\big(\sigma\,\delta A_{3}\big)-\int d\sigma\wedge\delta A_{3}$ (3.4)
-The boundary term is $\int d(\sigma\,\delta A_{3})$.
-We set it to zero by appropriate boundary conditions.
-The remaining term is $\int d\sigma\wedge\delta A_{3}$ and given that $\delta A_{3}$ is arbitrary in the interior, the Euler–Lagrange equation is,
-$d\sigma\Big(\frac{\Lambda}{\mu^{4}}\Big)=0$ (3.5)
-Now because $\sigma$ depends on spacetime only through $\Lambda$, this implies that $d(\Lambda/\mu^{4})=0$ provided $\sigma^{\prime}\neq 0$. Therefore $\Lambda$ is constant in spacetime.
-Repeating the same steps for $\hat{\sigma}(\lambda^{4}/\mu^{4})\int\hat{F}_{4}$, we have after varying with respect to $\hat{A}_{3}$,
-$d\hat{\sigma}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)=0$ (3.6)
-Assuming $\hat{\sigma}^{\prime}\neq 0$ we conclude that $d(\lambda^{4}/\mu^{4})=0$. Hence $\lambda$ is constant in spacetime and since $\lambda>0$ we can take $\lambda$ itself to be a global constant.
-Therefore, the four-form sectors force $\Lambda$ and $\lambda$ to be global constants. That is, they do not propagate and they are numbers fixed by boundary data and fluxes.
-2.) Varying the action with respect to the metric and identifying the local Einstein equations:
-We now compute the variation of each metric dependent term with respect to $g^{\mu\nu}$ where we will be using the identity $\delta\sqrt{-g}=-\dfrac{1}{2}\sqrt{-g}\,g_{\mu\nu}\,\delta g^{\mu\nu}$.
-The variation of the Einstein–Hilbert part gives the usual $M_{*}^{2}G_{\mu\nu}$ on the left hand side of the field equations.
-The variation of $T_{s}\int\sqrt{-g}$ gives,
-$\delta\big(-T_{s}\int\sqrt{-g}\big)=-T_{s}\int\delta\sqrt{-g}=\frac{T_{s}}{2}\int\sqrt{-g}\,g_{\mu\nu}\,\delta g^{\mu\nu}$ (3.7)
-This means the geometric tension contributes a stress tensor equal to $T_{s}g_{\mu\nu}$ with a minus sign when moved to the right hand side of Einstein equations.
-The variation of $\Lambda\int\sqrt{-g}$ is identical in structure and it gives a negative contribution $\Lambda g_{\mu\nu}$ to the right hand side of the field equations as well.
-Now we vary the matter action with respect to the metric, which reads,
-$S_{m}[g,\psi;\lambda]=\int\sqrt{-g}\,\,L_{m}(\lambda^{-2}\,g,\psi)$ (3.8)
-We define the physical matter stress tensor by the standard definition,
-$T^{(m)}_{\mu\nu}\equiv-\frac{2}{\sqrt{-g}}\,\frac{\delta S_{m}}{\delta g^{\mu\nu}}$ (3.9)
-The above definition ensures that the matter equations are consistent with diffeomorphism invariance and with the Bianchi identities. The full variation of the matter part yields,
-$\delta S_{m}=-\frac{1}{2}\int\sqrt{-g}\,T^{(m)}_{\mu\nu}\,\delta g^{\mu\nu}$ (3.10)
-Collecting all terms, the total variation with respect to $g^{\mu\nu}$ gives,
-$\delta S=\int\sqrt{-g}\,\Big[\frac{M_{*}^{2}}{2}\,G_{\mu\nu}+\frac{1}{2}\big(T_{s}+\Lambda\big)g_{\mu\nu}-\frac{1}{2}\,T^{(m)}_{\mu\nu}\Big]\,\delta g^{\mu\nu}$ (3.11)
-Stationarity for arbitrary $\delta g^{\mu\nu}$ gives,
-$\frac{M_{*}^{2}}{2}\,G_{\mu\nu}+\frac{1}{2}\big(T_{s}+\Lambda\big)g_{\mu\nu}-\frac{1}{2}\,T^{(m)}_{\mu\nu}=0$ (3.12)
-Multiplying by 2, the local Einstein field equations read,
-$M_{*}^{2}\,G_{\mu\nu}=T^{(m)}_{\mu\nu}-\big(T_{s}+\Lambda\big)\,g_{\mu\nu}$ (E)
-It is important to note that the four-form terms do not contribute to the metric equations.
-They are metric independent in the form we wrote, since $\int F_{4}$ and $\int\hat{F}_{4}$ do not carry $\sqrt{-g}$ or $g^{\mu\nu}$.
-Hence they add no local stress tensor.
-So, we now have the local field equations where we see that the geometric tension $T_{s}$ and the variable $\Lambda$ both appear as volume terms and therefore enter as $(\Lambda+T_{s})\,g_{\mu\nu}$.
-The matter sector contributes its usual stress tensor $T^{(m)}_{\mu\nu}$.
-3.) Varying the action with respect to $\Lambda$:
-Only two parts of the action (3.1) depend on $\Lambda$.
-The volume term $\Lambda\int\sqrt{-g}$ and the top-form term $\sigma(\Lambda/\mu^{4})\int F_{4}$.
-Therefore, computing the variation and treating $\Lambda$ as a global number, we find,
-$\delta S\big|_{\Lambda}=-\delta\Lambda\int d^{4}x\,\sqrt{-g}+\sigma^{\prime}\Big(\frac{\Lambda}{\mu^{4}}\Big)\,\frac{\delta\Lambda}{\mu^{4}}\,\int F_{4}$ (3.13)
-Setting the equation above this to zero for arbitrary $\delta\Lambda$, we obtain the first global constraint,
-$\int d^{4}x\,\sqrt{-g}=\frac{\sigma^{\prime}\left(\Lambda/\mu^{4}\right)}{\mu^{4}}\,\int F_{4}$ (C1)
-This equation tells us that the spacetime four-volume is proportional to the four-form flux weighted by $\sigma^{\prime}$.
-Note that it does not fix the local metric rather it only relates global numbers.
-With this, we have one global relation that involves $\Lambda$ but no local fields directly.
-4.) Varying the action with respect to $\lambda$:
-Again, only two parts of the action (3.1) depend on $\lambda$, which are the matter action $S_{m}$ and the top-form term $\hat{\sigma}(\lambda^{4}/\mu^{4})\int\hat{F}_{4}$.
-First consider the matter action which we compute $\delta S_{m}$ under a small variation of the global $\lambda$.
-$\delta S_{m}\bigg|_{\lambda}=\int\sqrt{-g}\,\frac{\partial L_{m}(\lambda^{-2}\,g,\,\psi)}{\partial(\lambda^{-2}g_{\mu\nu})}\,\delta(\lambda^{-2}g_{\mu\nu})$ (3.14)
-Let $h_{\mu\nu}\equiv\lambda^{-2}g_{\mu\nu}$ and its variation under a variation of $\lambda$ at fixed $g$ is,
-$\delta h_{\mu\nu}=\delta(\lambda^{-2})\,g_{\mu\nu}=-2\,\lambda^{-3}\,\delta\lambda\,g_{\mu\nu}$ (3.15)
-So we have,
-$\int\sqrt{-g}\;\frac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\;\delta(h_{\mu\nu})=-2\,\int\sqrt{-g}\;\lambda^{-3}\,\frac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\,g_{\mu\nu}\,\delta\lambda$ (3.16)
-In order to obtain $\dfrac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\,g_{\mu\nu}$ , we first vary $S_{m}$ with respect to the metric tensor $g_{\mu\nu}$, which gives,
-$\delta S_{m}=-\int\sqrt{-g}\;\lambda^{-2}\,\frac{\partial L_{m}(h_{\mu\nu},\psi)}{\partial(h_{\mu\nu})}g_{\alpha\mu}\,g_{\beta\nu}\delta g^{\mu\nu}$ (3.17)
-and use the definition of the matter stress energy tensor (3.9), which gives,
-$T_{\mu\nu}^{(m)}=2\,\lambda^{-2}\,\frac{\partial L_{m}(h_{\mu\nu},\psi)}{\partial(h_{\mu\nu})}g_{\alpha\mu}\,g_{\beta\nu}$ (3.18)
-Taking the trace, we find,
-$T^{(m)}=g^{\mu\nu}\,T_{\mu\nu}^{(m)}=2\,\lambda^{-2}\,\frac{\partial L_{m}(h_{\mu\nu},\psi)}{\partial(h_{\mu\nu})}g_{\mu\nu}$ (3.19)
-Solving for $\dfrac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\,g_{\mu\nu}$, we obtain,
-$\dfrac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}g_{\mu\nu}=\frac{\lambda^{2}\,T^{(m)}}{2}$ (3.20)
-Substituting into equation (3.16), we get,
-$\int\sqrt{-g}\;\frac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\;\delta(h_{\mu\nu})=-\frac{\delta\lambda}{\lambda}\int\sqrt{-g}\,T^{(m)}$ (3.21)
-Using (3.21) in (3.14), we obtain,
-$\delta S_{m}\bigg|_{\lambda}=-\frac{\delta\lambda}{\lambda}\int\sqrt{-g}\,T^{(m)}$ (3.22)
-Now we vary the top-form term $\hat{\sigma}(\lambda^{4}/\mu^{4})\int\hat{F}_{4}$, which gives,
-$\delta\Big[\hat{\sigma}\!\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\int\hat{F}_{4}\Big]=\hat{\sigma}^{\prime}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\,\frac{4\lambda^{3}\,\delta\lambda}{\mu^{4}}\,\int\hat{F}_{4}$ (3.23)
-Writing both pieces together, that is equations (3.21) and (3.23), the full variation of $S_{m}$ with respect to the global $\lambda$ then reads,
-$\frac{\delta\lambda}{\lambda}\,\int\sqrt{-g}\,T^{(m)}+\hat{\sigma}^{\prime}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\,\frac{4\lambda^{3}\,\delta\lambda}{\mu^{4}}\,\int\hat{F}_{4}=0$ (3.24)
-Canceling $\delta\lambda$ which is arbitrary but not zero and multiplying by $\lambda$, we obtain the second global constraint,
-$\int d^{4}x\,\sqrt{-g}\,T^{(m)}=\frac{4\,\lambda^{4}}{\mu^{4}}\,\hat{\sigma}^{\prime}\left(\frac{\lambda^{4}}{\mu^{4}}\right)\,\int\hat{F}_{4}$ (C2)
-We now have a second global relation that connects the spacetime integral of the matter trace to the flux of $\hat{F}_{4}$.
-The geometric tension $T_{s}$ does not appear because it does not depend on $\lambda$.
-5.) Eliminating fluxes to relate $\Lambda$ to the spacetime average of $T^{(m)}$:
-Equations (C1) and (C2) contain the fluxes $\int F_{4}$ and $\int\hat{F}_{4}$. These fluxes are constant numbers fixed by boundary conditions. We define the spacetime average of a quantity $X$ as
-$\langle X\rangle=\dfrac{\int\sqrt{-g}\,X}{\int\sqrt{-g}}$ (3.25)
-Combining (C1) and (C2) we may then write,
-$\langle T^{(m)}\rangle=4\,\lambda^{4}\,\frac{\hat{\sigma}^{\prime}(\lambda^{4}/\mu^{4})}{\sigma^{\prime}(\Lambda/\mu^{4})}\,\frac{\int\hat{F}_{4}}{\int F_{4}}$ (3.26)
-The right hand side is a pure number fixed by the values of the global variables and by the flux ratio. In the canonical sequestering normalization this relation is equivalent to,
-$\Lambda=\frac{1}{4}\,\langle T^{(m)}\rangle$ (3.27)
-This is the standard global identity in four-form sequestering and follows from the homogeneity in $\lambda$ together with the two Lagrange multiplier equations.
-Therefore, we have a global identity that expresses $\Lambda$ as one quarter of the spacetime average of the matter trace.
-6.) Performing the constant shift test and identify the invariant local combination:
-Consider a rigid shift of the matter Lagrangian by a constant $\Delta V$.
-Equivalently, we may replace $L_{m}\to L_{m}-\Delta V$ so that the added term in the action is $\int\sqrt{-g}\,\lambda^{4}\Delta V$.
-Using equation (3.9), we see that this simply changes the matter stress $T^{(m)}_{\mu\nu}$ and trace of matter stress $T^{(m)}$ by,
-$\Delta T^{(m)}_{\mu\nu}=-\Delta V\,g_{\mu\nu},\qquad\Delta T^{(m)}=-4\,\Delta V$ (3.28)
-The global relation then yields,
-$\Delta\Lambda=\frac{1}{4}\,\Delta\langle T^{(m)}\rangle=-\,\Delta V$ (3.29)
-Therefore the matter-vacuum local combination that appears in the Einstein equations is invariant,
-$T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}\;\to\;\big(T^{(m)}_{\mu\nu}-\Delta Vg_{\mu\nu}\big)-\big(\Lambda-\Delta V\big)g_{\mu\nu}=T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$ (3.30)
-All constant contributions from the matter sector cancel locally between $T^{(m)}_{\mu\nu}$ and $\Lambda g_{\mu\nu}$.
-This includes radiative corrections that enter as shifts of the vacuum energy in $L_{m}$.
-7.) Identifying the leftover geometric NG-term:
-Returning to the local Einstein equations (E),
-$M_{*}^{2}\,G_{\mu\nu}=T^{(m)}_{\mu\nu}-\big(T_{s}+\Lambda\big)\,g_{\mu\nu}$ (3.31)
-We rewrite the right hand side as,
-$\big[T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}\big]-T_{s}\,g_{\mu\nu}$ (3.32)
-The bracket is invariant under constant shifts of the matter Lagrangian.
-Therefore any constant vacuum energy from the matter sector is neutralized by the four-form enforced value of $\Lambda$.
-The term $T_{s}g_{\mu\nu}$ is untouched by this mechanism because $T_{s}$ does not depend on $\lambda$ and does not appear in the global trace equation.
-It is a purely geometric constant and it contributes a cosmological term to the Einstein equations that cannot be removed or adjusted by the four-form constraints.
-Therefore the effective cosmological term that remains in the local field equations is $T_{s}$.
-Thus, the neutralization does not act on $T_{s}$. The residual cosmological constant is $T_{s}$ as claimed.
-8.) General Remarks:
-A.
-Bianchi identities and conservation: Taking the divergence of both sides of the Einstein equations, the left hand side obeys $\nabla^{\mu}G_{\mu\nu}=0$ by the contracted Bianchi identity.
-The right hand side has $\nabla^{\mu}T^{(m)}_{\mu\nu}=0$ because the matter action is diffeomorphism invariant and couples minimally to $g_{\mu\nu}$.
-The term $(T_{s}+\Lambda)\,g_{\mu\nu}$ has zero divergence because $T_{s}$ and $\Lambda$ are global constants.
-Hence the equations are consistent with the Bianchi identities.
-B.
-Role of the top forms in the stress tensor: In the action (3.1) the four-form integrals carry no metric dependence and therefore do not contribute any local stress.
-This is why the local equations contain only $T^{(m)}_{\mu\nu}$ and the volume terms.
-If one used a different four-form model that included metric dependent contractions of $F_{\mu\nu\rho\sigma}$ with itself, then the four-form would carry local stress and would behave like a dynamical cosmological constant.
-That is the Brown–Teitelboim mechanism.
-Our conclusion is specific to the sequestering setup where the four-forms serve as global Lagrange multipliers rather than as local energy carriers.
-C.
-Radiative stability: Matter loop corrections shift the matter vacuum energy by constants of arbitrarily large size.
-These shifts are neutralized locally because the bracket $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$ is invariant.
-Gravitational loops can renormalize the geometric constant $T_{s}$ since it multiplies the volume element.
-In an effective field theory description one treats $T_{s}$ as a renormalized parameter that must be fixed by observation.
-So the key point is that matter loops do not destabilize the observed value once $T_{s}$ is chosen.
-We will discuss this point at length in the next Section.
-D.
-Scope of the theorem.
-The argument required only that the matter sector couples through a single global Weyl factor $\lambda$.
-It required that the geometric tension $T_{s}$ does not carry any explicit $\lambda$ dependence.
-If one added a non minimal coupling that made $T_{s}$ scale with $\lambda$, then the cancellation pattern would change and the result would not follow.
-Our setup excludes such couplings by assumption.
-9.) Einstein Equations for a pure matter vacuum:
-Consider a pure matter vacuum with $T^{(m)}_{\mu\nu}=-Vg_{\mu\nu}$ for some constant $V$.
-The trace is $T^{(m)}=-4\,V$ and the spacetime average of the trace equals $\langle T^{(m)}\rangle=-4V$.
-The global constraint gives $\Lambda=\dfrac{1}{4}\langle T^{(m)}\rangle=-V$.
-The matter-vacuum bracket therefore vanishes,
-$T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}]=0$ (3.33)
-and the Einstein equations reduce to,
-$M_{*}^{2}G_{\mu\nu}=-T_{s}\,g_{\mu\nu}$ (3.34)
-The geometric term $T_{s}$ in turn receives correction from graviton self loops and graviton-matter loops as we show in the subsequent Section.
-Therefore, it can be naturally identified with the residual observed Dark Energy.
-This conclusion is consistent with the general statement of the theorem and also makes the no geometric sequestering mechanism very transparent.
-$\square$
-4 Gravitational EFT, Graviton Loops and the Residual Term $T_{s}$
-Our goal is to present the low energy gravitational effective action in a way that keeps the discussion compact while preserving all structural information, see [36, 37].
-We use the background field method [39, 40], expand only the metric around a fixed background and keep matter fields unexpanded in this section.
-Background diffeomorphism invariance then forces the one particle irreducible functional $\Gamma$ to be a sum of background covariant operators.
-This locks in the tensor structure of counterterms and of the finite Wilsonian terms.
-Sequestering relation will then isolate the surviving constant source as the geometric volume tension $T_{s}$.
-We introduce a background metric $\bar{g}_{\mu\nu}$ and the fluctuation
-$g_{\mu\nu}\equiv\bar{g}_{\mu\nu}+\kappa\,h_{\mu\nu},\qquad\kappa^{2}\equiv 32\pi G\equiv\frac{4}{M_{*}^{2}}$ (4.1)
-The action is still given by (3.1),
-$S[g,\psi;\Lambda,\lambda]=\int d^{4}x\,\sqrt{-g}\,\Big(\tfrac{M_{*}^{2}}{2}\,R-T_{s}-\Lambda\Big)+S_{m}[\lambda^{-2}g,\psi]+S_{\text{4-forms}}[\Lambda,\lambda]$ (4.2)
-where the four form sector enforces the global constant constrains on $\Lambda$ and $\lambda$.
-We couple a source $J^{\mu\nu}$ to the metric fluctuation and define the gauge fixed partition function,
-$Z[\bar{g},J]=\int\mathcal{D}h\,\mathcal{D}c\,\mathcal{D}\bar{c}\,\mathcal{D}\psi\;e^{i\Big(S[\bar{g}+\kappa h,\psi]+S_{\text{gf}}[\bar{g},h]+S_{\text{gh}}[\bar{g},h,c,\bar{c}]+\int d^{4}x\,\sqrt{-\bar{g}}\,J^{\mu\nu}h_{\mu\nu}\Big)}$ (4.3)
-where $S_{\text{gf}}[\bar{g},h]$ and $S_{\text{gh}}[\bar{g},h,c,\bar{c}]$ are the gravitational gauge fixing and ghost Lagrangians respectively as discussed below.
-We now define the connected generator,
-$W[\bar{g},J]\equiv-\,i\ln Z[\bar{g},J]$ (4.4)
-The mean fluctuation field is therefore,
-$\hat{h}_{\mu\nu}(x)\equiv\frac{1}{\sqrt{-\bar{g}}}\frac{\delta W[\bar{g},J]}{\delta J^{\mu\nu}(x)}$ (4.5)
-We have introduced $\hat{h}$ in order to Legendre transform to the one-particle-irreducible functional.
-This produces vertex functions, clean Ward identities and the tadpole condition that fixes the background.
-The Legendre transform gives,
-$\Gamma[\bar{g},\hat{h}]\equiv W[\bar{g},J]-\int d^{4}x\,\sqrt{-\bar{g}}\,J^{\mu\nu}\hat{h}_{\mu\nu},\qquad\frac{1}{\sqrt{-\bar{g}}}\frac{\delta\Gamma}{\delta\hat{h}_{\mu\nu}}=-J^{\mu\nu}$ (4.6)
-Setting $J=0$ enforces the stationarity condition
-$\left.\frac{\delta\Gamma[\bar{g},\hat{h}]}{\delta\hat{h}_{\mu\nu}}\right|_{J=0}=0$ (4.7)
-By construction we choose the background so that $\hat{h}=0$ solves (4.7). The vacuum functional is then,
-$\Gamma[\bar{g}]\equiv\Gamma[\bar{g},\hat{h}]\big|_{\hat{h}=0}$ (4.8)
-This object generates the quantum-corrected field equations for the background metric.
-We now fix diffeomorphisms with a background covariant gauge.
-A standard choice is de Donder gauge built from $\bar{\nabla}$, where $\bar{\nabla}$ (the covariant derivative defined on background metric $\bar{g}$) is the unique torsion–free, metric–compatible connection satisfying,
-$\bar{\nabla}_{\rho}\bar{g}_{\mu\nu}=0,\qquad\bar{T}^{\rho}{}_{\mu\nu}\equiv\bar{\Gamma}^{\rho}{}_{\mu\nu}-\bar{\Gamma}^{\rho}{}_{\nu\mu}=0$ (4.9)
-and where,
-$\bar{\Gamma}^{\rho}{}_{\mu\nu}=\tfrac{1}{2}\,\bar{g}^{\rho\sigma}\big(\partial_{\mu}\bar{g}_{\nu\sigma}+\partial_{\nu}\bar{g}_{\mu\sigma}-\partial_{\sigma}\bar{g}_{\mu\nu}\big)$ (4.10)
-Using de Donder’s gauge, we have,
-$F_{\mu}^{\text{grav}}=\bar{\nabla}^{\nu}h_{\mu\nu}-\frac{1}{2}\bar{\nabla}_{\mu}h,\qquad\mathcal{L}_{\text{gf}}^{\text{grav}}=-\frac{1}{2\alpha_{g}}\,\sqrt{-\bar{g}}\,\bar{g}^{\mu\nu}F_{\mu}F_{\nu}$ (4.11)
-The Faddeev–Popov ghost action is then [41],
-$\Delta_{\ \nu}^{\mu}=\bar{\nabla}^{2}\delta^{\mu}_{\ \nu}+\bar{R}^{\mu}_{\ \nu},\qquad\mathcal{L}_{\text{gh}}^{\text{grav}}=-\sqrt{-\bar{g}}\,\bar{c}_{\mu}\,\Delta_{\ \nu}^{\mu}\,c^{\nu}$ (4.12)
-Let $\xi^{\mu}(x)$ generate an infinitesimal diffeomorphism.
-Background diffeomorphisms act as ordinary diffeomorphisms on $\bar{g}$ and as tensorial Lie derivatives on all quantum fields and on the source $J$,
-$\delta_{\text{bg}}\bar{g}_{\mu\nu}=\bar{\nabla}_{\mu}\xi_{\nu}+\bar{\nabla}_{\nu}\xi_{\mu},\qquad\delta_{\text{bg}}h_{\mu\nu}=\mathcal{L}_{\xi}h_{\mu\nu},$
-$\delta_{\text{bg}}c^{\mu}=\mathcal{L}_{\xi}c^{\mu},\qquad\delta_{\text{bg}}\bar{c}_{\mu}=\mathcal{L}_{\xi}\bar{c}_{\mu},\qquad\delta_{\text{bg}}J^{\mu\nu}=\mathcal{L}_{\xi}J^{\mu\nu}$ (4.13)
-Because $S_{\text{gf}}$ and $S_{\text{gh}}$ are built from $\bar{g}$ and $\bar{\nabla}$, they are invariant under $\delta_{\text{bg}}$.
-The classical action $S[\bar{g}+\kappa h,\psi]$ is diffeomorphism invariant by construction.
-Furthermore, we adopt a background-covariant regularization and measure, for instance dimensional regularization [38], so that $Z$ and $W$ are invariant when $J$ transforms as a background tensor.
-The Legendre transform then gives the Ward identity,
-$\delta_{\text{bg}}\Gamma[\bar{g},\hat{h}]=0,\qquad\delta_{\text{bg}}\hat{h}_{\mu\nu}=\mathcal{L}_{\xi}\hat{h}_{\mu\nu}$ (4.14)
-If we set $\hat{h}=0$, we obtain $\delta\,\Gamma[\bar{g}]=0$, which implies that $\Gamma[\bar{g}]$ is invariant under diffeomorphisms of $\bar{g}$ and therefore it can only be an integral of scalar densities built from $\bar{g}_{\mu\nu}$, the Riemann tensor of $\bar{g}$ and background covariant derivatives.
-This is the structural reason why the effective action organizes into background-covariant operators.
-Equation (4.14) is the reason why the effective action admits an expansion in background covariant operators.
-Every local counterterm and every local Wilsonian term (see [51, 42, 43, 44]) is an integral of $\sqrt{-\bar{g}}$ times a scalar built from $\bar{R}$, $\bar{R}_{\mu\nu}$, $\bar{R}_{\mu\nu\rho\sigma}$, and $\bar{\nabla}$ and it takes the universal form for the local part,
-$\Gamma^{\text{local}}[\bar{g}]=\int d^{4}x\,\sqrt{-\bar{g}}\,\Big[c_{0}(\mu)+\frac{1}{2}\,c_{2}(\mu)\,R+c_{4}^{(1)}(\mu)\,R^{2}+c_{4}^{(2)}(\mu)\,R_{\mu\nu}\,R^{\mu\nu}+\cdots\Big]$ (4.15)
-The non-local part contains structures such as $R\log(-\bar{\Box}/\mu^{2})R$ [49, 45, 50].
-These modify long distance kernels but do not act as constant background sources such as dark energy because they contain derivatives.
-In particular, the coefficient $c_{0}(\mu)$ renormalizes the volume term, that is $T_{s}$, the coefficient $c_{2}(\mu)$ renormalizes the Einstein–Hilbert coupling (Newton’s constant $G$), the $c_{4}^{(i)}(\mu)$ term renormalizes curvature squared operators and higher terms contain more derivatives and are suppressed at low curvature.
-The parameters $c_{n}(\mu)$ run with the renormalization scale $\mu$. We write,
-$\mu\,\frac{d}{d\mu}\,c_{0}(\mu)=\beta_{0}(\mu),\qquad\mu\,\frac{d}{d\mu}\,c_{2}(\mu)=\beta_{2}(\mu),\qquad\mu\,\frac{d}{d\mu}\,c^{(i)}_{4}(\mu)=\beta^{(i)}_{4}(\mu),\quad\text{etc.}$ (4.16)
-The precise values depend on the field content and on the gauge fixing and on the regulator.
-We now combine tree level and loop corrected coefficients that multiply the same operators. We define renormalized parameters at the scale $\mu$ by
-$T_{s}(\mu)$ $\equiv T_{s}^{\text{bare}}-c_{0}(\mu),$
-$M_{*}^{2}(\mu)$ $\equiv M_{*\,\text{bare}}^{2}+c_{2}(\mu),$
-$\alpha_{1,2}(\mu)$ $\equiv c_{4}^{(1,2)}(\mu)$ (4.17)
-The local gravitational part of the effective action can therefore be written as,
-$\Gamma^{\text{grav}}_{\text{local}}=\int d^{4}x\,\sqrt{-g}\,\Big(-T_{s}(\mu)+\frac{M_{*}^{2}(\mu)}{2}R+\alpha_{1}(\mu)R^{2}+\alpha_{2}(\mu)R_{\mu\nu}R^{\mu\nu}+\cdots\Big)$ (4.18)
-Varying and adding the matter part gives the local field equations,
-$M_{*}^{2}(\mu)\,G_{\mu\nu}=[T^{(m)}_{\mu\nu}-\Lambda\,g_{\mu\nu}]-T_{s}(\mu)\,g_{\mu\nu}+\mathcal{H}_{\mu\nu}[R^{2},\dots]$ (4.19)
-where $\mathcal{H}_{\mu\nu}$ collects higher derivative tensors from curvature squared terms.
-These tensors are suppressed at low curvature and do not act like a constant source.
-The combination $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$ is invariant under rigid shifts $L_{m}\to L_{m}-\Delta V$ in the sequestering setup, hence constant vacuum energy inside $S_{m}$ cancels out of the local equations.
-So in the field equation (4.19), the net constant vacuum source after sequestering is
-$\rho_{\rm vac}(\mu)\,g_{\mu\nu}=-T_{s}(\mu)\,g_{\mu\nu}$ (4.20)
-At late times one can match this constant source to the observed dark energy density,
-$\rho_{\Lambda}^{\rm obs}=T_{s}(\mu_{\rm IR})$ (4.21)
-where $\mu_{\rm IR}$ is an infrared matching scale appropriate for cosmology. This identification fixes one linear combination of low energy parameters using data.
+The construction developed so far captures only the uniform tension sector of the intrinsic membrane.
+However, once the membrane of space is allowed to support an internal hidden $U(1)$ gauge field, the Nambu–Goto description is no longer the most complete local nonlinear effective action.
+The natural generalization is then of Dirac–Born–Infeld type.
+In the limit that the hidden gauge field is switched off, the DBI action reduces to the Nambu–Goto volume term.
+When the gauge field is present, it resums the nonlinear response of the membrane to finite worldvolume fluxes and provides a finite-field completion of the Maxwell regime.
+We therefore turn next to the DBI completion of the intrinsic membrane picture and ask what kind of hidden magnetic sectors it admits.
+2.4 The DBI Completion of the Elastic Space Picture
+A convenient starting point is the four-dimensional DBI gauge action (we work in units where $\hbar=c=1$),
+$S_{\rm DBI}=-T_{s}\int d^{4}x\,\sqrt{-\det\left(g_{\mu\nu}+\lambda_{h}F_{\mu\nu}\right)}\,.$ (2.31)
+Factoring out the metric determinant gives,
+$S_{\rm DBI}=-T_{s}\int d^{4}x\,\sqrt{-g}\,\sqrt{\det\left(\delta^{\mu}{}_{\nu}+\lambda_{h}F^{\mu}{}_{\nu}\right)}\,.$ (2.32)
+Since $F_{\mu\nu}$ is antisymmetric, the determinant identity in four dimensions yields,
+$\det\left(\delta^{\mu}{}_{\nu}+\lambda_{h}F^{\mu}{}_{\nu}\right)=1+\frac{\lambda_{h}^{2}}{2}F_{\mu\nu}\,F^{\mu\nu}-\frac{\lambda_{h}^{4}}{16}(F_{\mu\nu}\tilde{F}^{\mu\nu})^{2}\,.$ (2.33)
+Therefore,
+$S_{\rm DBI}=-T_{s}\int d^{4}x\,\sqrt{-g}\,\sqrt{1+\frac{\lambda_{h}^{2}}{2}F_{\mu\nu}\,F^{\mu\nu}-\frac{\lambda_{h}^{4}}{16}(F_{\mu\nu}\tilde{F}^{\mu\nu})^{2}}\,.$ (2.34)
+Writing $\lambda_{h}=\beta_{h}^{-1}$, one obtains,
+$S_{\rm DBI}=-T_{s}\int d^{4}x\,\sqrt{-g}\,\sqrt{1+\frac{1}{2\beta_{h}^{2}}\,F_{\mu\nu}\,F^{\mu\nu}-\frac{1}{16\beta_{h}^{4}}\,(F_{\mu\nu}\tilde{F}^{\mu\nu})^{2}}\,.$ (2.35)
+Note that the equation above in weak fields reduces to,
+$S_{\rm DBI}=-T_{s}\int d^{4}x\,\sqrt{-g}\,\left[1+\frac{1}{4\,\beta_{h}^{2}}\,F_{\mu\nu}\,F^{\mu\nu}\right]\,.$ (2.36)
+When $\beta_{h}^{2}=T_{s}$ as in the Maxwell normalization, we have,
+$S_{\rm DBI}=-\int d^{4}x\,\sqrt{-g}\,\left[T_{s}+\frac{1}{4}\,F_{\mu\nu}\,F^{\mu\nu}\right]\,$, (2.37)
+with the Lagrangian density,
+$\mathcal{L}_{\rm DBI}=-T_{s}-\frac{1}{4}F_{\mu\nu}\,F^{\mu\nu}\,.$ (2.38)
+At this point, the minimal physical content of the construction becomes clear.
+Once space is treated as an effective DBI membrane, the theory already contains two basic ingredients.
+The first is the vacuum-like or cosmological contribution $T_{s}$, which in the present framework is identified with the residual observed vacuum sector.
+The second is an abelian gauge sector arising from the DBI structure itself, whose weak-field limit is the $U(1)$ field appearing above.
+In this work, we regard that $U(1)$ as hidden.
+The reason is simple.
+Nothing in the bare effective action requires it to be identified with ordinary electromagnetism, and in particular the visible matter content and coupling assignments do not follow automatically from the minimal DBI sector alone.
+One may certainly imagine more elaborate ultraviolet completions in which Standard Model gauge fields and matter are engineered from intersecting or stacked brane configurations, but such constructions are highly model dependent and often rather contrived.
+Our purpose here is more modest.
+We only use the minimal DBI completion as an intrinsic effective description of space, which naturally furnishes a vacuum term together with a hidden $U(1)$ sector, without assuming that this sector is the electromagnetic field of the Standard Model.
+3 Late-Time Running of the Tension of Space
+We now consider a late-time microscopic origin for the running of the space tension.
+The purpose of the present section is not to provide a complete ultraviolet model of hidden-sector interaction.
+Rather, we construct a proof of concept effective description in which the intrinsic hidden gauge sector already suggested by the DBI completion of space develops, at late times, a defect-like reservoir that can exchange energy with the vacuum-like tension sector.
+We consider here an additional hidden-sector transition at late times, whose detailed microphysics we do not attempt to derive here, but whose coarse-grained cosmological effects can still be studied consistently.
+3.1 Hidden $U(1)\to\mathbb{Z}_{n}$ breaking and Flux-Tube Defects
+We now extend the intrinsic hidden $U(1)$ gauge sector of space, already introduced in the DBI description, by a complex scalar field $\Phi$ carrying charge $n$ under that same hidden Abelian symmetry.
+At the level of a minimal effective description, the hidden sector is modeled by extending the DBI Lagrangian of space, equation (2.38), so as to include a hidden Higgs field.
+The corresponding effective Lagrangian is
+$\mathcal{L}_{\rm hid}=-\frac{1}{4}Z(\sigma)\,F_{\mu\nu}\,F^{\mu\nu}+|D_{\mu}\Phi|^{2}-\frac{\lambda_{\Phi}}{4}\left(|\Phi|^{2}-v_{\Phi}^{2}(\sigma)\right)^{2}\,$, (3.1)
+with,
+$D_{\mu}\Phi=\left(\partial_{\mu}-i\,ne_{h}A_{\mu}\right)\Phi\,.$ (3.2)
+Here $A_{\mu}$ is the hidden gauge field, $e_{h}$ is the hidden gauge coupling, and $\sigma$ is a slow collective variable that will be used below to parameterize the dynamical part of the space-tension sector.
+The functions $Z(\sigma)$ and $v_{\Phi}(\sigma)$ allow the hidden sector to communicate with the tension sector at the effective level.
+When the hidden Higgs field condenses, at a late enough time in cosmic history (as discussed in Sec. 4.2),
+$\langle\Phi\rangle=\frac{v_{\Phi}}{\sqrt{2}}\neq 0\,$, (3.3)
+the continuous hidden gauge symmetry is not necessarily destroyed completely.
+Because $\Phi$ carries charge $n$, the vacuum remains invariant under the discrete subgroup $\mathbb{Z}_{n}\subset U(1)_{h}$ [60, 61].
+At the effective level, the late-time symmetry-breaking pattern is therefore,
+$U(1)_{h}\longrightarrow\mathbb{Z}_{n}\,.$ (3.4)
+Owing to this condensation, the late-time hidden sector enters a topologically nontrivial broken phase in which localized flux-tube defects can form, see [62, 63, 64, 65].
+The basic mechanism is standard.
+During the transition, different causal regions need not select the same phase of the hidden complex scalar $\Phi$.
+As a result, the phase of $\Phi$ can acquire a nontrivial winding around certain closed loops in space.
+Such a winding cannot be removed continuously while remaining everywhere in the broken phase.
+At least along some line-like locus, the scalar magnitude must therefore be driven back toward zero so that the phase becomes ill-defined there.
+These line-like loci are precisely the cores of the hidden flux tubes.
+The gauge field then adjusts so as to maintain finite energy away from the core.
+In particular, the covariant gradient energy of the Higgs field requires the phase winding of $\Phi$ to be compensated by an azimuthal gauge configuration, so that the hidden magnetic flux becomes trapped inside the defect core rather than remaining as a freely propagating long-range field.
+Equivalently, once the hidden $U(1)$ sector is Higgsed to a residual discrete subgroup, the gauge boson acquires a mass in the broken phase, and the bulk expels magnetic flux in the usual Meissner-like manner [62, 66, 67].
+The flux is therefore energetically forced into narrow, string-like tubes.
+For a winding number $k$, the trapped flux is quantized, schematically as,
+$\Phi_{B}=\oint A_{\theta}\,r\,d\theta=\frac{2\pi k}{ne_{h}}\,,\qquad k\in\mathbb{Z}\,$, (3.5)
+so that the resulting objects are stable localized carriers of hidden magnetic flux.
+In this way, the late-time symmetry breaking does not merely leave behind a bath of free hidden magnetic radiation.
+Rather, it reorganizes the relevant hidden energy into an extended network of flux-tube defects whose microscopic width is set by the inverse hidden Higgs and gauge masses, while their lengths may be macroscopic on cosmological scales.
+The detailed nonequilibrium dynamics of the transition, including defect formation, intercommutation, loop production, and the subsequent network evolution [64, 65], are not required at the phenomenological level.
+For the present proof-of-concept analysis, it is sufficient to treat the magnetic flux network in coarse-grained form as a hidden flux-tube reservoir with an effective energy density $\rho_{\rm flux}$ and an effective equation-of-state parameter $w_{\rm flux}$.
+Once such a reservoir is present, the next question is how it communicates with the vacuum-like sector associated with the intrinsic tension of space.
+In particular, if the hidden defect network can exchange energy with the dark-energy sector, then the total tension can no longer remain purely rigid at late times.
+This motivates a decomposition of the space tension into a constant background piece and a dynamical contribution, from which a time-dependent effective equation of state for dark energy naturally emerges.
+3.2 Splitting the Tension Sector and the Effective Equation of State of Dark Energy
+If we assume that the energy carried by the magnetic flux tube defects produced due to the breaking of $U(1)$ gauge symmetry by the hidden Higgs field can be exchanged with the vacuum sector, then the spatial tension $T_{s}$ identified in Sec.
+2 splits into a constant background piece and a dynamical transferred piece,
+$T(t)=T_{s}+\delta T_{s}(t)\,,\qquad\dot{T}_{s}=0\,.$ (3.6)
+Here $T_{s}$ denotes as before the rigid background contribution that survives after the ultraviolet physics has played its role.
+It is identified with the dominant constant part of the observed dark-energy scale as mentioned earlier.
+By contrast, $\delta T_{s}$ represents the late-time correction induced by energy exchange with the hidden flux-tube sector.
+Only $\delta T_{s}$ is allowed to participate in the exchange.
+This is physically important as the constant background $T_{s}$ is the baseline vacuum sector itself and is understood as a built-in background parameter inherited from hard ultraviolet complete physics.
+In that sense it is rigid and cannot be treated as a reservoir that either drains energy into other sectors or receives energy back from them.
+It is simply the irreducible vacuum baseline that remains after the ultraviolet problem has already been resolved.
+What can evolve is only an additional contribution deposited on top of this baseline by later hidden-sector dynamics.
+Therefore $\delta T_{s}$ denotes the excess energy temporarily stored in the tension sector.
+Any subsequent relaxation, unloading, or transfer should then be understood as acting only on $\delta T_{s}$, not on the underlying vacuum piece $T_{s}$ itself.
+The late-time dynamics does not remove the intrinsic vacuum tension of space.
+It only perturbs the total tension away from its rigid baseline and may later allow that excess contribution to decay.
+At the background level, we continue to identify the dark-energy density with the total space tension,
+$\rho_{\rm DE}(t)\;\equiv\;T(t)\;=\;T_{s}+\delta T_{s}(t)\,.$ (3.7)
+Since $\rho_{\rm DE}$ changes with time, or equivalently the scale factor $a$, its effective equation of state is also time dependent and reads (see Appendix A),
+$w_{\rm eff}(a)=-1-\frac{1}{3}\frac{d\ln T}{d\ln a}=-1-\frac{1}{3\bigl(T_{s}+\delta T_{s}\bigr)}\frac{d\delta T_{s}}{d\ln a}\,.$ (3.8)
+One can therefore see that if $\delta T_{s}$ changes with the scale factor, then $w_{\rm eff}$ changes as well, giving a dynamical equation of state for dark energy.
+Equation (3.8) makes the central point of the present framework explicit.
+Any late-time exchange that modifies $\delta T_{s}$ immediately induces a departure of $w_{\rm eff}$ from 1-1, even though the rigid background contribution $T_{s}$ itself remains untouched.
+In this sense, the running of dark energy is not attributed to the erosion of the vacuum baseline, but to the evolution of an additional tension component deposited on top of it by hidden-sector dynamics.
+What remains is to represent this exchange in a simple local effective description.
+This is the purpose of the next subsection.
+3.3 Pedagogical Lagrangian for the coupled Tension and Flux Sectors
+The splitting introduced above is sufficient to define the background dark-energy sector and its effective equation of state, but it does not yet specify a local field-theoretic channel through which the hidden flux reservoir communicates with the dynamical part of the tension.
+To make that interaction more concrete, it is useful to introduce a simple pedagogical effective action in which the transferred part of the tension is encoded by a slowly varying scalar degree of freedom [35, 68].
+This provides a compact local framework for organizing the couplings between the tension sector, the hidden gauge field, and the hidden Higgs field.
+To write a local effective action for a dynamical exchange, it is convenient to represent the transferred part of the tension sector by a slowly varying scalar degree of freedom $\sigma$ such that,
+$T(\sigma)=T_{s}+\delta T_{s}(a,\sigma)\,.$ (3.9)
+A simple effective action follows from augmenting equation (3.1) with a term to encode the dynamics of the $\sigma$ sector and replace $T_{s}$ with $T=T_{s}+\delta T_{s}$,
+$S=\int d^{4}x\,\sqrt{-g}\,\Bigg[-\Bigl(T_{s}$ $+\delta T_{s}(a,\sigma)\Bigr)-\frac{1}{4}Z(\sigma)\,F_{\mu\nu}\,F^{\mu\nu}$ (3.10)
+$+|D_{\mu}\Phi|^{2}-\frac{\lambda_{\Phi}}{4}\left(|\Phi|^{2}-v_{\Phi}^{2}(\sigma)\right)^{2}-\frac{K_{\sigma}}{2}\,(\partial^{\mu}\sigma\,\partial_{\mu}\sigma)\Bigg]\,$,
+where the quantity $K_{\sigma}$ is a constant.
+Furthermore, $\delta T_{s}(\sigma)$ makes explicit that only the dynamical correction to the tension is carried by $\sigma$, while $T_{s}$ remains constant.
+The couplings $Z(\sigma)$ and $v_{\Phi}(\sigma)$ provide an effective communication channel between the hidden flux-tube sector and the tension sector.
+As mentioned earlier, the present treatment is intentionally phenomenological.
+We do not attempt to solve the full defect-formation and string-network dynamics from the field theory above.
+Instead, we move to a coarse-grained background description.
+3.4 Coarse-grained Continuity Equations
+Let $\rho_{\rm flux}$ denote the coarse-grained energy density of the hidden flux-tube network, and let its effective pressure be parameterized by,
+$p_{\rm flux}=w_{\rm flux}\,\rho_{\rm flux}\,.$ (3.11)
+For an almost frozen string-like reservoir one expects $w_{\rm flux}$ to be negative, while for a more rapidly evolving network it may be closer to zero.
+In the present proof-of-concept treatment we leave $w_{\rm flux}$ as an effective parameter.
+We now assume that the transferred part of the tension sector and the flux-tube reservoir exchange energy through a bidirectional channel.
+The corresponding continuity equations are taken to be (see Appendix B),
+$\dot{\rho}_{\rm flux}+3\,H\bigl(1+w_{\rm flux}\bigr)\rho_{\rm flux}$ $=-\,Q\,$, (3.12)
+$\dot{\delta T_{s}}$ $=Q\,.$ (3.13)
+where $Q$ is the amount of energy density exchanged per unit time. The exchange term is modeled phenomenologically as,
+$Q=\Gamma H\left(\rho_{\rm flux}-\kappa\,T_{s}\right)\,,\qquad\Gamma\,>\,0\,,\qquad 0\,\leq\,\kappa\,\leq\,1\,.$ (3.14)
+where $\Gamma$ is a dimensionless coupling parameter.
+This form captures the simplest balanced communication channel between the two dynamical sectors.
+It should be mentioned here that the phenomenological exchange law introduced above was written in its minimal form by taking the coupling parameter $\Gamma$ to be constant.
+However, once the transferred part of the tension sector is encoded by the scalar degree of freedom $\sigma$ as discussed in Sec.
+3.3, it is natural to expect that the efficiency of the exchange channel itself may depend on the state of that sector.
+This expectation is reinforced by the fact that the same scalar already controls the couplings $Z(\sigma)$ and $v_{\Phi}(\sigma)$ in the local effective action.
+However, here we keep the analysis simple and treat $\Gamma$ as a constant.
+When the hidden flux-tube reservoir dominates over the stored dynamical correction to the space tension, that is when
+$\rho_{\rm flux}\,>\,\kappa\,T_{s}\,$, (3.15)
+one has Q> 0Q\,>\,0, so energy flows from the hidden defect sector into the space tension. By contrast, when
+$\rho_{\rm flux}\,<\,\kappa\,T_{s}\,$, (3.16)
+one has Q< 0Q\,<\,0, so the excess dynamical part of the tension sector relaxes back into the hidden flux sector.
+Importantly, the constant background piece $T_{s}$ never participates in this exchange.
+Substituting (3.14) into (3.12) and (3.13), the coupled background system becomes
+$\dot{\rho}_{\rm flux}+3H\bigl(1+w_{\rm flux}\bigr)\rho_{\rm flux}$ $=-\,\Gamma H\left(\rho_{\rm flux}-\kappa\,T_{s}\right)\,$, (3.17)
+$\dot{\delta T_{s}}$ $=\Gamma H\left(\rho_{\rm flux}-\kappa\,T_{s}\right)\,.$ (3.18)
+If we further assume that the energy density of the flux-tube-network is some fraction $\eta$ of the residual vacuum energy $T_{s}$, then,
+$\rho_{\rm flux}=\eta(t)\,T_{s}\,,\qquad 0\,\leq\,\eta(t)\,\leq\,1$ (3.19)
+which gives,
+$\dot{\eta}+3H\bigl(1+w_{\rm flux}\bigr)\eta$ $=-\Gamma H(\eta-\kappa)\,$, (3.20)
+$\dot{\delta T_{s}}$ $=\Gamma HT_{s}(\eta-\kappa)\,.$ (3.21)
+Note that here we have assumed $\eta(t)$ to be a time dependent function because the energy density of the flux-tube-network will decrease as a function of time due to cosmic expansion.
+Writing the evolution in terms of the e-folding variable,
+$a=e^{N}\,,\qquad N\equiv\ln a\,,\qquad\frac{d}{dN}=\frac{1}{H}\frac{d}{dt}\,$, (3.22)
+the system becomes,
+$\frac{d\eta}{dN}+3\bigl(1+w_{\rm flux}\bigr)\eta$ $=-\Gamma(\eta-\kappa)$ (3.23)
+$\frac{d\,\delta T_{s}}{dN}$ $=\Gamma\,T_{s}(\eta-\kappa)\,.$ (3.24)
+Equation (3.23) further gives,
+$\frac{d\eta}{dN}=-\Bigl[3\bigl(1+w_{\rm flux}\bigr)+\Gamma\Bigr]\eta+\Gamma\,\kappa\,$, (3.25)
+If $w_{\rm flux}$, $\Gamma$, and $\kappa$ are treated as constants, the solution for $\eta(a)$ is,
+$\eta(a)=\eta_{\infty}+\bigl(\eta_{i}-\eta_{\infty}\bigr)\left(\frac{a}{a_{i}}\right)^{-m}\,$, (3.26)
+with,
+$m=3\bigl(1+w_{\rm flux}\bigr)+\Gamma\,,\qquad\eta_{\infty}=\frac{\Gamma\,\kappa}{m}\,.$ (3.27)
+In equation (3.26), the term $a_{i}$ and $\eta_{i}$ denote the initial value of both, the scale factor and the hidden flux fraction, at that point in time when the hidden Higgs field condenses, thereby breaking the symmetry of the DBI gauge field from $U(1)\to\mathbb{Z}_{n}$.
+We further write,
+$m\;=\;3\bigl(1+w_{\rm flux}\bigr)+\Gamma\;=\;n+\Gamma\,,\qquad n\equiv 3\bigl(1+w_{\rm flux}\bigr)$ (3.28)
+Now, using equation (3.24) in (3.8), we get,
+$w_{\rm eff}(a)=-1-\frac{\Gamma\,T_{s}}{3\bigl[T_{s}+\delta T_{s}(a)\bigr]}\bigl(\eta(a)-\kappa\bigr)\,.$ (3.29)
+This makes the phantom-divide crossing transparent. If the hidden flux-tube sector is initially large enough such that,
+$\eta(a)>\kappa\,$, (3.30)
+then,
+$\frac{d\,\delta T_{s}}{dN}>0\qquad\Longrightarrow\qquad w_{\rm eff}(a)<-1\,.$ (3.31)
+If,
+$\eta(a)=\kappa\,$, (3.32)
+then,
+$\frac{d\,\delta T_{s}}{dN}=0\qquad\Longrightarrow\qquad w_{\rm eff}(a)=-1\,.$ (3.33)
+At later times, the hidden defect reservoir redshifts and depletes. Once the balance reverses, one has,
+$\eta(a)<\kappa\,$, (3.34)
+then,
+$\frac{d\,\delta T_{s}}{dN}<0\qquad\Longrightarrow\qquad w_{\rm eff}(a)>-1\,.$ (3.35)
+The model therefore admits, at the level of a simple coarse-grained effective description, a transient crossing of the phantom divide.
+In this way the hidden flux-tube reservoir generated by the late-time $U(1)_{h}\to\mathbb{Z}_{n}$ transition provides a physically transparent proof of concept for a mild low-redshift running of the effective dark-energy sector.
+The coarse-grained system derived above yields an explicit late-time evolution for $\eta(a)$, $\delta T_{s}(a)$, and hence for the effective dark-energy equation of state $w_{\rm eff}(a)$.
+The natural next step is to ask whether this predicted running is phenomenologically relevant.
+Rather than confronting the full cosmological likelihood directly, we adopt a simpler benchmark strategy and compare the model evolution with the effective CPL description preferred by current late-time data.
+This allows us to assess, in a transparent way, how closely the present running-tension scenario can approach the observationally favored region in the $(w_{0},w_{a})$ plane.
+4 A DESI-CPL Benchmark and the Flux-Tube Running-Tension Scenario
+The coarse-grained system developed in Section 3 yields a late-time prediction for the hidden flux fraction $\eta(a)$, the dynamical tension shift $\delta T_{s}(a)$, and therefore the effective dark-energy equation of state $w_{\rm eff}(a)$.
+The purpose of the present section is to test this prediction against a simple observational benchmark suggested by current late-time data.
+More specifically, we ask whether the flux-tube driven running of the tension sector can reproduce the kind of mild low-redshift evolution commonly summarized in the Chevallier–Polarski–Linder, or CPL plane.
+Our aim here is diagnostic rather than inferential.
+We do not attempt a full end-to-end cosmological likelihood analysis of the hidden-sector model.
+Instead, we use a compressed benchmark built from the public $w_{0}w_{a}$ posterior and compare the resulting CPL target with the explicit curve $w_{\rm eff}(z)$ predicted by the running-tension model.
+This is sufficient for the proof-of-concept purpose of the present work, which is to determine whether a late hidden defect reservoir can induce a phenomenologically relevant departure from a strictly constant dark-energy sector.
+A conceptually similar compressed-benchmark comparison, applied to a different dynamical dark-energy model based on spatial phonons, was also adopted in [55].
+4.1 Benchmark Data and Effective CPL Target
+[…]
+Let,
+$\hat{\theta}=\begin{pmatrix}\hat{w}_{0}\\ \hat{w}_{a}\end{pmatrix}$ (4.1)
+denote the weighted posterior mean of the released chains, and let $C$ denote the corresponding weighted covariance matrix.
+The compressed benchmark used in the present section is then completely specified by the pair $(\hat{\theta},C)$.
+The Gaussian compression of the released posterior samples, namely the extraction of the weighted mean vector $\hat{\theta}$ and covariance matrix $C$, was performed with a custom Python script, w_0_w_a_python.py, which is available at [73].
+Using the released chains, we obtain,
+$\hat{\theta}=\begin{pmatrix}-0.828209\\ -0.744750\end{pmatrix},\qquad C=\begin{pmatrix}0.004112&-0.016695\\ -0.016695&0.085227\end{pmatrix}$. (4.2)
+Equation (4.2) is the only observational input required for the analysis that follows.
+The benchmark is therefore intentionally lightweight.
+It does not retain non-Gaussian structure in the full posterior, nor does it track correlations with the other cosmological parameters that would enter a full global fit.
+That loss of information is acceptable for the limited purpose of the present section, which is only to assess whether the flux-tube mechanism can generate a low-redshift evolution with roughly the right shape and amplitude.
+The corresponding CPL trajectory is [71, 72],
+$w_{\rm CPL}(z)=w_{0}+w_{a}\frac{z}{1+z}.$ (4.3)
+Substituting the benchmark mean values from equation (4.2) gives the specific target curve against which the running-tension model will be tested.
+Although the observational benchmark is expressed in the CPL plane, the quantity predicted directly by the present model is not the pair $(w_{0},w_{a})$, but the full redshift-dependent equation of state $w_{\rm eff}(a)$ generated by the running tension $T(a)=T_{s}+\delta T_{s}(a)$.
+Accordingly, the comparison is first performed at the level of the equation-of-state curve itself.
+We then project the model prediction $w_{\rm eff}(z)$, as given in equation (3.8), onto the CPL form of equation (4.3) in order to extract an effective pair $(w_{0}^{\rm mod},w_{a}^{\rm mod})$ for comparison with the benchmark values $(\hat{w}_{0},\hat{w}_{a})$.
+The projection is carried out over the interval $0\leq z\leq 1.6$ by means of an ordinary least-squares fit onto the basis $\{1,\,z/(1+z)\}$, evaluated on a dense redshift grid.
+4.2 Pointwise Fit of the Flux-Tube Running Tension Model to the DESI-CPL Curve
+We now test the flux-tube running-tension ansatz directly against the DESI-motivated effective CPL benchmark. At the background level, we take the dark-energy density to be the total space tension,
+$\rho_{\rm DE}(a)=T(a)=T_{s}+\delta T_{s}(a)\,$, (4.4)
+and we compare the resulting effective equation of state $w_{\rm eff}(a)$ to the compressed DESI-CPL target,
+$\hat{w}_{\rm CPL}(a)=\hat{w}_{0}+\hat{w}_{a}(1-a)\,,\qquad(\hat{w}_{0},\hat{w}_{a})=(-0.828209,-0.744750)\,.$ (4.5)
+The comparison is performed over the interval $0\leq z\leq 1.6$ which corresponds to,
+$a_{i}\in[a_{\min},1]\,,\qquad a_{\min}=\frac{1}{1+1.6}\simeq 0.384615\,.$ (4.6)
+For $a<a_{i}$, we set $w_{\rm eff}(a)=-1$, while for $a\geq a_{i}$ we use the analytic solution of the coarse-grained exchange equations discussed in Sec.
+3.4 with the initial condition $\delta T_{s}(a_{i})=0$.
+The scan is carried out under the constraints,
+$n=3\left(1+w_{\rm flux}\right)\in\{2,3,4\}\,,\qquad 0\leq\eta_{i}\leq 1\,,\qquad 0\leq\kappa\leq 1\,,\qquad\Gamma\geq 0\,.$ (4.7)
+To quantify the agreement, we sample the interval $a\in[a_{\min},1]$ with 100 uniformly spaced points $a_{j}$ and minimize the pointwise root mean square error,
+${\rm RMSE}=\sqrt{\frac{1}{100}\sum_{j=1}^{100}\left[w_{\rm eff}(a_{j})-\hat{w}_{\rm CPL}(a_{j})\right]^{2}}\,.$ (4.8)
+For each parameter point we also project the model curve back onto CPL form by an ordinary least squares fit on the basis $\{1,z/(z+1)\}$, thereby obtaining an effective pair $(w_{0}^{\rm mod},w_{a}^{\rm mod})$.
+The coarse scan used the uniform grid,
+$N_{a_{i}}=31\,,\qquad N_{\eta_{i}}=31\,,\qquad N_{\Gamma}=61\,,\qquad N_{\kappa}=31\,$, (4.9)
+which corresponds to,
+$31\times 31\times 61\times 31=1{,}815{,}771$ (4.10)
+models for each fixed $n$, or 5,447,3135{,}447{,}313 models in total for the restricted set $n=2,3,4$.
+The Python script performing this scan is available as scan_flux_fit.py at [73].
+Within this restricted family, the best fit is obtained for $n=2$ which is summarized in Table 1 below.
+$n$ $a_{i}$ $\eta_{i}$ $\Gamma$ $\kappa$ RMSE $w_{0}^{\rm mod}$ $w_{a}^{\rm mod}$
+2 0.384615 1.0000 1.45 0.1667 0.07471 -0.91231 -0.45474
+Table 1: Best coarse-grid pointwise fit of the flux-tube running-tension model to the DESI-motivated effective CPL benchmark.
+Performing a local refinement around $n=2$, fixing $a_{i}=a_{\min}$ and scanning more densely over,
+$\eta_{i}\in[0.90,1.00]\,,\qquad\Gamma\in[1.20,1.70]\,,\qquad\kappa\in[0.05,0.30]\,$, (4.11)
+using a $41\times 101\times 101$ grid. This yields the refined best-fit point,
+$n=2\,,\qquad a_{i}=0.384615\,,\qquad\eta_{i}=1.0000\,,\qquad\Gamma=1.45\,,\qquad\kappa=0.1525\,$, (4.12)
+with,
+${\rm RMSE}=0.07463\,.$ (4.13)
+Projecting the model curve onto the basis $\{1,z/(1+z)\}$, we obtain,
+$(w_{0}^{\rm mod},w_{a}^{\rm mod})=(-0.91529,-0.45716)\,.$ (4.14)
+Two qualitative features are noteworthy.
+First, the fit drives the transition epoch to the earliest value allowed within the DESI leverage range, namely $a_{i}=a_{\min}\approx 0.384$.
+The preferred onset of the transition therefore lies not in the early universe, but in the relatively recent cosmological past, well after recombination and deep into the late-time matter-dominated era.
+In the language of the present model, this means that the hidden Higgs field condensation is favored to occur as a genuinely late-time event, precisely in the redshift interval where a mild evolution of the dark-energy sector could be phenomenologically relevant.
+At the same time, it is important to emphasize that the late onset of the hidden Higgs condensation should not be viewed as implausible merely because it occurs at low redshift.
+Since the relevant sector is hidden, its symmetry-breaking history is not tied to that of the visible Standard Model and can therefore be considerably more flexible.
+In particular, delayed condensation may be realized in several ways, for example through a temperature-dependent hidden effective potential, through couplings to a slowly varying background degree of freedom such as the tension-sector scalar, or through radiative and threshold effects that postpone the onset of symmetry breaking until late times.
+We do not attempt to build such a mechanism explicitly in the present work.
+Nevertheless, the hidden nature of the sector makes a late-time condensation a reasonable possibility rather than an exceptional one, and a more detailed realization of this transition would be an important direction for future work.
+Second, it also drives the initial hidden flux fraction to its maximal value, $\eta_{i}=1$.
+This indicates that, within the minimal single-reservoir ansatz, the model prefers both an early onset and a maximally populated hidden defect sector in order to mimic the DESI-motivated low-redshift running.
+At the same time, the resulting effective CPL projection remains visibly displaced from the benchmark target, although the mismatch is not so large as to preclude further improvement.
+This should be viewed as encouraging rather than discouraging.
+The present construction is only a proof of concept based on a highly simplified coarse-grained description, and there remains substantial room for improvement once the hidden-sector dynamics, the defect-network evolution, and the exchange channel with the tension sector are modeled in greater detail.
+It is therefore plausible that a more complete realization of the same basic idea could reduce, or potentially close, the remaining gap to the observational benchmark.
+Furthermore, substituting $n=2$ into equation (3.28) gives,
+$w_{\rm flux}=\frac{2}{3}-1=-\frac{1}{3}\,.$ (4.15)
+Therefore, in our model, the best-fit hidden flux sector behaves like an effectively string-like network of line-like defects.
+This is physically consistent with the underlying picture developed in Section 3, where the late-time broken phase reorganizes the hidden magnetic energy into flux tubes rather than into a bath of freely propagating relativistic quanta.
+A component with $w_{\rm flux}=-1/3$ dilutes more slowly than radiation and sits precisely at the boundary between decelerating and curvature-like behavior, which makes it cosmologically more persistent at late times.
+In the present framework, this is important because it allows the hidden defect reservoir to survive long enough, and with sufficient strength, to exchange energy with the dynamical part of the space tension and thereby induce a nontrivial evolution of $w_{\rm eff}(a)$.
+Given below is a table and a plot comparing the equation of state of dark energy as obtained from DESI and flux-tube running tension model, both projected to a CPL basis.
+$a$ $z$ $w_{\rm CPL}^{\rm DESI}$ $w_{\rm eff}^{\rm model}$ $\Delta w$
+0.384615 1.600000 -1.286517 -1.409625 -0.123108
+0.459207 1.177665 -1.230964 -1.175142   0.055823
+0.533800 0.873362 -1.175412 -1.084349   0.091063
+0.608392 0.643678 -1.119859 -1.040081   0.079778
+0.682984 0.464164 -1.064307 -1.015535   0.048772
+0.757576 0.320000 -1.008754 -1.000711   0.008043
+0.832168 0.201681 -0.953202 -0.991182 -0.037980
+0.906760 0.102828 -0.897650 -0.984749 -0.087100
+1.000000 0.000000 -0.828209 -0.979315 -0.151106
+Table 2: Comparison between the DESI-motivated CPL benchmark and the best-fit flux-tube running-tension model at representative points across the interval $0\leq z\leq 1.6$.
+The model values are computed from the running-tension relation $w_{\rm eff}(a)=-1-\dfrac{1}{3}\dfrac{d\ln T}{d\ln a}$, with $T(a)=T_{s}+\delta T_{s}(a)$, using the best-fit restricted-$n$ parameters $n=2$, $a_{i}=1/2.6$, $\eta_{i}=1$, $\Gamma=1.45$, and $\kappa=0.1525$.
+Here $\Delta w\equiv w_{\rm eff}^{\rm model}-w_{\rm CPL}^{\rm DESI}$.
+Figure 1: Comparison between the DESI-motivated CPL benchmark equation of state and the best-fit flux-tube running-tension model over the range $0\leq z\leq 1.6$.
+The blue curve shows the CPL form $w_{\rm CPL}(a)=w_{0}+w_{a}(1-a)$ with the benchmark parameters $(w_{0},w_{a})=(-0.828209,-0.744750)$, while the orange curve shows the model prediction obtained from the running-tension relation $w_{\rm eff}(a)=-1-\dfrac{1}{3}\dfrac{d\ln T}{d\ln a}$, with $T(a)=T_{s}+\delta T_{s}(a)$.
+The model curve is evaluated using the best-fit restricted-$n$ point from the 100-point scan, namely $n=2$, $a_{i}=1/2.6$, $\eta_{i}=1$, $\Gamma=1.45$, and $\kappa=0.1525$.
+The figure illustrates that, although the model captures a qualitatively evolving low-redshift equation of state, it does not reproduce the DESI-CPL curve pointwise across the full fitting interval.
+For additional perspective, it is useful to compare the model-implied CPL pair obtained here with the two analyses presented by Chudaykin and Kunz [74].
+In their baseline $w_{0}w_{a}$CDM analysis of the CMB + DESI + SN combination, they found,
+$w_{0}=-0.821\pm 0.065,\qquad w_{a}=-0.76^{+0.31}_{-0.26},$ (4.16)
+which is very close to our quoted value in (4.2), obtained from a rather simple analysis of the cobaya chains. On the other hand, their modified-gravity Horndeski analysis yielded,
+$w_{0}^{H}=-0.856\pm 0.062,\qquad w_{a}^{H}=-0.53^{+0.28}_{-0.26}.$ (4.17)
+The effective CPL pair inferred from the present running-tension construction (equation (4.14)) lies in the same broad region of the $(w_{0}^{H},w_{a}^{H})$ plane.
+This proximity indicates that the phenomenological proof-of-concept developed in the present work captures a nontrivial part of the same late-time trend identified in more comprehensive studies.
+In this sense, the hidden-sector flux-tube mechanism should be viewed as a physically transparent toy realization of evolving dark energy, rather than as a mere numerical coincidence.
 5 Discussion
-We developed a minimal framework in which space is modeled as an elastic three membrane that fills spacetime.
-The dynamics are captured by an intrinsic Nambu–Goto volume term $S_{s}=-T_{s}\int\sqrt{-g}\,d^{4}x$.
-This fixes the geometric stress tensor to $T^{(s)}_{\mu\nu}=-T_{s}\,g_{\mu\nu}$ and identifies the membrane tension with a uniform vacuum energy of space.
-We then embedded this term into a matter vacuum sequester framework that employs global four form constraints to cancel constant contributions from the matter sector.
-The variation shows that the global variable adjusts to neutralize the matter vacuum energy while a purely geometric unit operator (the operator which behaves like vacuum energy) carries no dependence on that variable and is not in the cancellation channel.
-As a result the local field equations retain a residual contribution $T_{s}\,g_{\mu\nu}$.
-We interpret this as the effective cosmological term.
-Subsequently we performed the path integral over the gravitational fluctuation $h_{\mu\nu}$ and over the associated Faddeev–Popov ghosts only.
-We did not integrate over matter fluctuations so as to not “recompute” the pure matter loops.
-Their effect is assumed to have already been absorbed into the renormalized matter functional $S_{m}[\lambda^{-2}g,\psi]$ at the matching scale $\mu$.
-The outcome was a background diffeomorphism invariant one particle irreducible action $\Gamma[\bar{g}]$ with the structure shown in equation (4.15).
-The local part is an infinite covariant tower built from the curvature and its derivatives.
-The non-local part contains form factors that never reduce to a pure volume term and so do not influence vacuum energy.
-The geometric volume tension $T_{s}(\mu)$ arises from the renormalized coefficient of the unit operator and enters the field equations.
-This separation makes clear which ingredients can shift $T_{s}(\mu)$ and which ingredients only change higher derivative responses of the geometry.
-If one wishes to include matter-graviton mixed loops explicitly, one can extend the path integral to the matter fields fluctuations as well.
-The background field method and general covariance then enforce the same covariant operator basis that appears in equation (4.15).
-Mixed matter-graviton diagrams that contain at least one internal graviton line are local in the ultraviolet.
-They therefore renormalize the same coefficients that multiply curvature invariants as in equation (4.15).
-No new local scalar of lower canonical mass dimension exists beyond the unit operator and the Ricci scalar term.
-As a result the only change is in the numerical values of the Wilson coefficients.
-Concretely, mixed matter–graviton loops shift the same $c_{0}(\mu)$ and therefore in turn shift $T_{s}(\mu)\equiv T_{s}^{\rm bare}-c_{0}(\mu)$.
-They also shift $c_{2}(\mu)$ which renormalizes the constant $G$ along with shifting the coefficients of curvature squared invariants and beyond.
-The tensorial form of each contribution is fixed by background diffeomorphisms and does not depend on the details of the gauge fixing.
-It can further be arranged so that pure matter functional $S_{m}[\lambda^{-2}g,\psi]$ will still continue to encode matter self interactions and pure matter loops at the chosen matching scale.
-The combined result is a single covariant effective action with renormalized coefficients $c_{0}(\mu),c_{1}(\mu),\cdots$ that capture both gravitational and mixed contributions in a uniform way.
-Therefore, the generalization to mixed matter-graviton graphs is straightfoward.
-With this in mind, one can then conveniently regard $T_{s}(\mu)$ as the fully renormalized geometric tension that already includes contributions from pure graviton self loops and from mixed matter–graviton loops because as mentioned before, extending the path integral from metric fluctuations to metric plus matter fluctuations only shifts the Wilson coefficients in the same covariant operator basis.
-The definition of $T_{s}(\mu)$ as the renormalized coefficient of the unit operator therefore captures all such shifts in a single parameter.
-One might try to redefine $\Lambda^{\prime}\equiv\Lambda-T_{s}$ so that the sequestering constraint cancels both the matter vacuum energy and the geometric volume term.
-This identification is physically not possible because in the standard sequestering framework which we are considering the global variables and four-forms are wired so that the variational constraint removes constant shifts of the matter Lagrangian, while the purely gravitational sector is inert under that constraint.
-An independent geometric unit operator $T_{s}\int\sqrt{-g}$ is therefore not neutralized unless one explicitly extends the mechanism to couple the same global sector to that operator.
-To also cancel $T_{s}$, one would need a different theory, one which couples a new rigid variable to a geometric density.
-This represents a different sequestering scheme and whereas we are working in the minimal framework which requires just two global Lagrange multipliers $\lambda$ and $\Lambda$.
-One can also see why such an identification is not possible in the minimal case because the operator $T_{s}\int\sqrt{-g}$ is blind to $\lambda$ and therefore cannot contribute to the $\lambda$ constrain.
-So it is correct to keep $\Lambda$ as the variable that enforces the global matter constraint and we keep $T_{s}(\mu)$ as an intrinsic geometric coupling that lives in the gravitational effective action.
-In the minimal sequester framework, the added advantage of having an un-sequestered geometric term is that after fixing covariant operator basis using the background field method, the term $T_{s}(\mu)$ automatically emerges as the renormalized coefficient of the unit operator after graviton and mixed matter–graviton loops are included.
-Nonlocal form factors appear but they do not produce a pure volume term and therefore do not modify this identification.
-The constant source in the field equations can therefore naturally be identified as $\rho_{\rm cav}(\mu)=T_{s}(\mu)$.
-At an infrared scale $\mu_{\rm IR}$ appropriate for cosmology one can then match $\rho_{\rm vac}(\mu_{\rm IR})$ to the observed dark energy density.
-This retains the virtue of sequestering for matter vacuum energy while leaving a natural geometric volume term that can account for the residual dark energy.
-It also avoids promoting geometric couplings to global multipliers and keeps conservation and locality manifest through the Bianchi identity.
+The present construction should be viewed as a proof of concept for a particular physical idea, namely that the late-time dark-energy sector may be interpreted as the intrinsic tension of space itself, and that mild departures from exact vacuum behavior may arise if a hidden defect reservoir exchanges energy with the dynamical part of that tension.
+In this framework, the observed cosmological term is not replaced by an ad hoc dark-energy fluid.
+Rather, one begins from an intrinsic membrane description of space, identifies its uniform tension with the residual vacuum sector, and then asks whether additional hidden-sector structure can perturb that vacuum contribution away from exact constancy at late times.
+A central outcome of the analysis is that the model does generate a nontrivial effective equation of state for dark energy.
+In particular, once the total tension is split into a rigid background piece and a dynamical correction, the exchange of energy between the hidden flux-tube reservoir and the latter induces a time-dependent $w_{\rm eff}(a)$.
+At the coarse-grained level, this setup also permits a transient crossing of the phantom divide.
+In this sense, the model succeeds in establishing the basic logical point that a running tension of space can mimic a mild low-redshift evolution of the dark-energy sector without requiring the vacuum baseline itself to be removed or destabilized.
+The benchmark fit discussed in Section 4 also offers a useful piece of physical guidance.
+Within the restricted family considered there, the preferred solution is driven toward the earliest transition epoch allowed by the DESI leverage range and toward the largest possible initial hidden flux fraction.
+Moreover, the best restricted fit corresponds to $n=2$, which implies $w_{\rm flux}=-1/3$.
+This is not radiation-like behavior.
+Rather, it is the equation of state expected of an effectively string-like network of line-like defects.
+That result is physically consistent with the picture developed in Section 3, where the late-time broken phase traps hidden magnetic flux into tube-like configurations instead of leaving it in a bath of freely propagating relativistic quanta.
+The fit therefore points back toward the underlying interpretation of the model, namely that a sufficiently persistent defect reservoir is favored if one wishes to generate an observable running of the tension sector.
+At the same time, the present treatment has important limitations.
+The first and most obvious one is that the hidden flux-tube network has not been modeled in detail.
+The late-time $U(1)_{h}\to\mathbb{Z}_{n}$ transition is introduced only at the effective level, and the subsequent defect formation and full network evolution are not computed microscopically.
+Instead, all of that complicated nonequilibrium dynamics is compressed into the effective quantities $\rho_{\rm flux}$ and $w_{\rm flux}$.
+This simplification is appropriate for a first proof-of-concept study, but it also means that the cosmological sector of the model is not yet tied to a fully derived defect-network history.
+A second limitation is that the exchange term between the hidden reservoir and the dynamical part of the tension sector is phenomenological.
+The form adopted in this work is simple, physically interpretable, and sufficient to illustrate the desired effect, but it has not been derived from a first-principles treatment of the local hidden-sector action.
+A more complete analysis would need to show how the effective transfer law emerges from the microscopic dynamics of the hidden Higgs field, the hidden gauge sector, and the evolving defect network.
+Closely related to this is the fact that the late-time phase transition itself is assumed rather than derived from a detailed cosmological history.
+In particular, the present work does not yet explain why the hidden-sector transition should occur precisely in the redshift interval where a DESI-like signal would be most relevant.
+A third limitation concerns the observational comparison.
+The analysis of Section 4 does not fit the full BAO, supernova, and CMB likelihoods directly.
+Instead, it compares the model against a compressed Gaussian description of the effective CPL posterior in the $(w_{0},w_{a})$ plane.
+This makes the comparison transparent and reproducible, and it is entirely adequate for a first diagnostic test of the idea.
+However, it also means that the resulting agreement should not be interpreted as a precision cosmological fit.
+The residual gap between the best-fit model curve and the benchmark target is therefore not surprising.
+Indeed, that mismatch should be read as a sign that the minimal single-reservoir ansatz is still too restrictive, not as evidence that the underlying physical idea must fail.
+There is substantial room for improvement once the hidden-sector dynamics, the network evolution, and the exchange mechanism are treated in greater detail.
+There is also an important conceptual limitation in scope.
+The present model does not solve the cosmological constant problem itself.
+The quantity $T_{s}$ is interpreted throughout as the small residual vacuum sector that remains after whatever ultraviolet physics is responsible for suppressing the enormous microscopic vacuum contributions has already done its job.
+In other words, the present framework is not a first-principles derivation of why the vacuum energy is tiny.
+Rather, it is an attempt to understand whether the surviving late-time vacuum sector may admit a natural effective description in terms of an intrinsic space tension, and whether hidden-sector dynamics can generate a mild departure from exact constancy around that already-small baseline.
+Finally, the present work is restricted to the homogeneous background evolution.
+It does not yet address perturbations in the hidden sector, the clustering properties of the defect reservoir, possible stability issues beyond the background level, or additional signatures that may arise outside the effective CPL plane.
+These questions are essential if the framework is to be developed into a more complete phenomenological model.
+Taken together, these limitations define a clear program for future work.
+The most immediate next step is to replace the coarse-grained reservoir description by a more explicit treatment of the hidden defect network and its late-time cosmological evolution.
+A second step would be to derive, or at least better motivate, the exchange channel from the underlying local action.
+A third would be to extend the analysis from the compressed CPL benchmark to a full likelihood-based confrontation with cosmological data.
+It is also worth noting that the tension-of-space framework appears flexible enough to support more than one late-time dynamical realization.
+In particular, the companion phonon model of Ref.
+[55] implements the same broad geometric viewpoint through a viscous longitudinal phonon sector and was found to closely track a DESI-motivated CPL benchmark over the DESI-sensitive redshift range.
 6 Conclusion
-We have modeled space as an elastic three membrane and identified its uniform volume tension with a geometric unit operator that contributes a stress proportional to the metric.
-The Nambu–Goto term $S_{s}=-T_{s}\int\sqrt{-g}\,d^{4}x$ fixes $T^{(s)}_{\mu\nu}=-T_{s}g_{\mu\nu}$ and allows a direct comparison with vacuum energy.
-Embedding this structure into a four–form sequester framework, we showed that global constraints neutralize constant matter contributions inside the invariant combination $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$, while the geometric unit operator does not couple to the global variable and therefore cannot be removed.
-The local field equations then take the renormalized effective form,
-$M_{*}^{2}(\mu)\,G_{\mu\nu}=\bigl[T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}\bigr]-T_{s}(\mu)\,g_{\mu\nu}+H_{\mu\nu}[R^{2},\nabla R,\ldots]$
-with higher–derivative tensors suppressed at low curvature.
-Using the background field method we argued that graviton loops and mixed matter–graviton loops renormalize the unit operator, so $T_{s}(\mu)$ runs with the scale.
-The residual constant source may be identified with dark energy at an infrared matching scale relevant for cosmology.
-An immediate next step is to extend the elastic space picture by introducing bulk modulus and bulk viscosity to space.
-In a homogeneous expansion a nonzero bulk viscosity produces an effective negative pressure that can evolve with the Hubble rate.
-This offers a natural route to time dependent dark energy and can help account for DESI indications of evolving dark energy.
-We will develop this viscous elastic space framework next.
+In this work, we have explored the possibility that the late-time dark-energy sector may be interpreted as the intrinsic tension of space itself.
+Starting from an intrinsic membrane description, we showed that a uniform space tension contributes to the gravitational field equations with exactly the tensor structure of vacuum energy.
+We then considered the Dirac–Born–Infeld completion of this picture, which naturally introduces a hidden Abelian gauge sector in addition to the vacuum-like tension term.
+[…]
+We then compared the resulting evolution with a compressed observational benchmark in the effective CPL plane.
+Although the minimal single-reservoir realization does not exactly reproduce the benchmark target, it is able to generate a nontrivial late-time running that lies in the phenomenologically relevant direction.
+In particular, the preferred restricted fit points toward an early onset of the transition within the DESI-sensitive range and toward a substantial hidden defect population, which is consistent with the basic physical picture developed in the model.
+The present framework should be interpreted as a proof of concept rather than as a complete theory of dark energy.
+We have not attempted to derive the late-time hidden-sector transition from first principles, nor have we modeled the full formation and evolution of the hidden flux-tube network.
+Likewise, the exchange law between the defect reservoir and the tension sector has been treated phenomenologically, and the observational comparison has been carried out at the level of a compressed CPL benchmark rather than through a full likelihood analysis.
+Even so, the construction demonstrates a logically consistent and physically suggestive possibility, in that once the vacuum sector is reinterpreted as an intrinsic tension of space, hidden-sector defect physics can naturally induce a mild running of the effective dark-energy component.
+There are several natural directions for future work.
+A more complete treatment should derive the effective transfer mechanism from the underlying hidden-sector dynamics, model the defect-network evolution in greater detail, and extend the present background-level study to perturbations and full cosmological likelihood tests.
+It would also be interesting to investigate whether modest extensions of the present setup can reduce the remaining gap to the observational benchmark.
+For these reasons, we regard the present model as a minimal starting point for a broader program in which the dark-energy sector is viewed not as an independent fluid, but as a dynamical manifestation of the intrinsic tension of space.
 7 Acknowledgment
-The author acknowledges the use of ChatGPT-5 (August 2025 version) to edit and refine tone and clarity in selected sections of this manuscript.
+The author acknowledges the use of ChatGPT-5.4 to edit and refine tone and clarity in selected sections of this manuscript.
 The model was used only to polish and rephrase text supplied by the author.
 All LLM assisted passages were reviewed and edited by the author, exercising the final editorial judgment.
+Appendix A Derivation of the effective equation of state
+We now derive the relation,
+$w_{\rm eff}(a)=-1-\frac{1}{3}\frac{d\ln T}{d\ln a}\,.$ (A.1)
+The starting point is the continuity equation for a homogeneous effective fluid in an FLRW background,
+$\dot{\rho}_{\rm DE}+3H\bigl(\rho_{\rm DE}+p_{\rm DE}\bigr)=0\,$, (A.2)
+where $H=\dot{a}/a$ is the Hubble parameter. Writing the pressure in the usual form,
+$p_{\rm DE}=w_{\rm eff}\,\rho_{\rm DE}\,$, (A.3)
+one finds,
+$\dot{\rho}_{\rm DE}+3H\bigl(1+w_{\rm eff}\bigr)\,\rho_{\rm DE}=0\,.$ (A.4)
+Solving for $w_{\rm eff}$ gives,
+$w_{\rm eff}=-1-\frac{\dot{\rho}_{\rm DE}}{3H\rho_{\rm DE}}\,.$ (A.5)
+Using,
+$\frac{d\ln\rho_{\rm DE}}{d\ln a}=\frac{1}{\rho_{\rm DE}}\,\frac{d\rho_{\rm DE}}{d\ln a}=\frac{\dot{\rho}_{\rm DE}}{H\rho_{\rm DE}}\,$, (A.6)
+Equation (A.5) therefore becomes,
+$w_{\rm eff}(a)=-1-\frac{1}{3}\frac{d\ln\rho_{\rm DE}}{d\ln a}.$ (A.7)
+In the present framework, the effective dark-energy density is identified with the tension of space,
+$\rho_{\rm DE}(a)\equiv T(a)\,.$ (A.8)
+Therefore,
+$w_{\rm eff}(a)=-1-\frac{1}{3}\frac{d\ln T}{d\ln a}\,.$ (A.9)
+If we further split the tension as,
+$T(a)=T_{s}+\delta T_{s}(a)\,,\qquad\dot{T}_{s}=0\,$, (A.10)
+then,
+$\frac{d\ln T}{d\ln a}=\frac{1}{T_{s}+\delta T_{s}}\frac{d\delta T_{s}}{d\ln a}\,.$ (A.11)
+Substituting this into Eq. (A.9), one obtains,
+$w_{\rm eff}(a)=-1-\frac{1}{3\bigl(T_{s}+\delta T_{s}\bigr)}\frac{d\delta T_{s}}{d\ln a}\,.$ (A.12)
+Appendix B Derivation of the coupled continuity equations
+We model the late-time interacting sector in coarse-grained form as the sum of two components, fist of which is a hidden flux-tube reservoir and second being the dynamical part of the space-tension sector.
+Accordingly, we split the total stress tensor as,
+$T^{\mu\nu}_{\rm tot}=T^{\mu\nu}_{\rm flux}+T^{\mu\nu}_{\delta T_{s}}\,.$ (B.1)
+We assume that the total interacting sector is covariantly conserved,
+$\nabla_{\mu}T^{\mu\nu}_{\rm tot}=0\,$, (B.2)
+but that the two individual components may exchange energy through an interaction current $Q^{\nu}$ according to,
+$\nabla_{\mu}T^{\mu\nu}_{\rm flux}=-Q^{\nu}\,,\qquad\nabla_{\mu}T^{\mu\nu}_{\delta T_{s}}=+Q^{\nu}\,.$ (B.3)
+For a homogeneous FLRW background, the transfer is taken to be purely temporal in the comoving frame,
+$Q^{\nu}=Q\,u^{\nu}\,$, (B.4)
+where $u^{\nu}$ is the background four-velocity and $Q$ is the scalar energy-transfer rate.
+The flux-tube sector is described as an effective fluid with,
+$p_{\rm flux}=w_{\rm flux}\,\rho_{\rm flux}\,.$ (B.5)
+Its continuity equation then becomes,
+$\dot{\rho}_{\rm flux}+3H\bigl(\rho_{\rm flux}+p_{\rm flux}\bigr)=-\,Q\,$, (B.6)
+or equivalently,
+$\dot{\rho}_{\rm flux}+3H\bigl(1+w_{\rm flux}\bigr)\rho_{\rm flux}=-\,Q\,.$ (B.7)
+We now treat the dynamical transferred part of the space tension as a vacuum-like component with,
+$\rho_{\delta T_{s}}=\delta T_{s}\,,\qquad p_{\delta T_{s}}=-\,\delta T_{s}\,.$ (B.8)
+Its continuity equation is therefore,
+$\dot{\rho}_{\delta T_{s}}+3H\bigl(\rho_{\delta T_{s}}+p_{\delta T_{s}}\bigr)=Q\,.$ (B.9)
+Since,
+$\rho_{\delta T_{s}}+p_{\delta T_{s}}=0\,$, (B.10)
+this reduces immediately to,
+$\dot{\delta T_{s}}=Q\,.$ (B.11)
+We thus obtain the coupled background system,
+$\dot{\rho}_{\rm flux}+3H\bigl(1+w_{\rm flux}\bigr)\rho_{\rm flux}$ $=-\,Q\,$, (B.12)
+$\dot{\delta T_{s}}$ $=Q\,.$ (B.13)
+These equations simply express that the hidden flux-tube sector loses exactly the amount of energy gained by the dynamical part of the space-tension sector, and conversely when the sign of $Q$ reverses.